### v1

# Spectral representations of interpolation spaces of reproducing kernel Hilbert spaces

Abstract
In statistical learning theory, interpolation spaces of the form $\mathrm{L}^{2},H]_{\theta,r}$, where $H$ is a reproducing kernel Hilbert space, are in widespread use. So far, however, they are only well understood for fine index $r=2$. We generalise existing results from $r=2$ to all possible values of $r$. In particular, we present a spectral decomposition of such spaces, analyse their embedding properties, and describe connections to the theory of Banach spaces of functions. We additionally present exemplary applications of our results to regularisation error estimation in statistical learning.
1 Introduction
Since their introduction, reproducing kernel Hilbert spaces (RKHS) have turned out as fundamental objects in various branches of mathematical theory as well as in applications. As they combine the geometry of a Hilbert space with the continuous evaluation functionals of a function space, they posses many desirable properties. Apart from applications in functional analysis and numerics, statistical learning theory takes advantage of this situation by considering a large variety of kernel-based algorithms.
Recent years showed that in the statistical analysis of such learning algorithms, certain interpolation spaces play a key role. To be more precise, for an RKHS $H$, spaces of the form
$[\mathrm{L}^{2}(\nu),H]_{\theta,2}$ (1)
are important for the analysis of kernel-based least squares regression, see e.g. [5], for estimations of the conditional mean embedding, see e.g. [11], and the analysis of learning curves, see e.g. [10]. Similarly, these spaces occur naturally for certain questions related to Gaussian processes, see e.g. [15], [9].
The main reason for choosing $2$ as the fine index in (1) is the spectral representation of these interpolation spaces. More precisely, the spaces (1) are given as the range of powers of the integral operator associated to the RKHS [18, Section 4]. Using eigenvalues and eigenfunctions of this operator hence yields many insights into interpolation spaces and turns out as very helpful in statistical learning theory.
Aside from these practical advantages of fine index $2$, considering other values turned out as beneficial. In particular, the following two cases are of special interest.
  • Fine index $1$ is of interest as it is allows for certain embedding that are not possible for other fine indexes, see e.g. [17, p. 4]
  • Fine index $\infty$ is of interest as it allows for a precise analysis of approximation errors, see e.g. [14, Theorem 3.1].
For these reasons, we consider real interpolation spaces of the general form
$[\mathrm{L}^{2}(\nu),H]_{\theta,r}\,\text{,}$ (2)
that is, we allow for all possible fine indexes $1\leq r\leq\infty$. Our aim is to present methods and results that treat such interpolation spaces for all fine indexes $r$ in a joint fashion. The results, we will derive address the following main questions:
  • Do the spaces (2) and their norms still possess a spectral representation?
  • The interpolation spaces consist of equivalence classes of functions at first. When can we identify the spaces (2) with Banach spaces consisting of functions?
  • When are the spaces (2) continuously embedded into $\mathrm{L}^{\infty}$?
The first question is clearly motivated by the practicality of spectral representations, as indicated by the papers mentioned earlier. Positive answers to the second question have applications in all situations where a point evaluation of functions is necessary. A main example is given by the theory of Gaussian processes in the context of machine learning, as paths cannot be equivalence classes but need to be defined everywhere, see e.g. [12, Assumption 1]. The third question will turn out to be tightly connected to the second one, but applications go beyond that. In particular, the statistical analysis of kernel-based learning algorithms relies on such embeddings, see e.g. [16, Theorem 7.23] [5, Section 3] [17, p. 4].
This work is organised as follows: We start by introducing the necessary notation and fundamental definitions in Section 2. Then the main results are provided in Section 3. We present an exemplary application to error estimation in statistical learning theory in Section 4. All proofs and the ideas behind the results will be given in Section 5. Finally, some auxiliary results can be found in two appendices.
2 Preliminaries
Given normed $\mathbb{R}$-vector spaces $E_{1}$ and $E_{2}$, we use $E_{1}\simeq E_{2}$ to denote isometric isomorphy, $E_{1}\triangleq E_{2}$ for norm equivalent spaces and $E_{1}\hookrightarrow E_{2}$ if $E_{1}\subset E_{2}$ and the embedding is continuous.
For a measurable space $(X,\mathcal{A})$, we denote the set of measurable functions $X\rightarrow\mathbb{R}$ by $\mathcal{L}^{0}(X)$. The set of bounded functions in $\mathcal{L}^{0}(X)$ is denoted by $\mathcal{L}^{\infty}(X)$ and this space is equipped with the supremum norm.
Let us assume that additionally, a measure $\nu$ is given on $X$. Then for $1\leq p<\infty$, we consider the space $\mathcal{L}^{p}(\nu)$ of all functions in $\mathcal{L}^{0}(X)$ with finite seminorm $\|f\|_{\mathcal{L}^{p}(\nu)}:=(\int_{X}|f|^{p}\mathrm{d}\nu)^{\frac{1}{p}}$. Similarly, $\mathcal{L}^{\infty}(\nu)$ denotes the space of essentially bounded functions in $\mathcal{L}^{0}(X)$ equipped with the seminorm $\|f\|_{\mathcal{L}^{\infty}(\nu)}:=\operatorname*{\mathrm{ess-sup}}_{x\in X}|f(x)|$. Furthermore, we consider equivalence classes $f]_{\sim}$ of functions that are $\nu$-almost everywhere equal to $f\in\mathcal{L}^{0}(X)$. For $E\subset\mathcal{L}^{0}(X)$, we denote the set of all equivalence classes of functions in $E$ as $E]_{\sim}$. In particular, for $1\leq p\leq\infty$, $\mathrm{L}^{p}(\nu):=[\mathcal{L}^{p}(\nu)]_{\sim}$ is equipped with the classical $\mathrm{L}^{p}$-norm inherited from $\mathcal{L}^{p}(\nu)$. These spaces are Banach spaces. For such indices $p$, we extend the typical arithmetic by $\frac{1}{\infty}:=0$. If $\nu$ is fixed and if $w:X\mapsto[0,\infty)$ is a measurable “weight” function, we denote the $\mathrm{L}^{p}$-space with respect to the measure $wd\nu$ by $\mathrm{L}^{p}(w)$. For the counting measure $\nu$, the same applies to $\ell^{p}(w)$.
Interpolation spaces
Let us introduce some background from the theory of interpolation spaces of the real interpolation method. For more details, see [3, Chapter 5] or [4, Chapter 3]. In the following, we only consider interpolation spaces of Banach spaces $E_{1}$ and $E_{2}$ where $E_{1}\hookrightarrow E_{2}$.
To define the real interpolation spaces, we use the so-called $K$-method. To this end, for all $x\in E_{2}$ and $t>0$, we define the $K$-functional as
$K(x,t,E_{2},E_{1}):=\inf_{y\in E_{1}}\left(\|x-y\|_{E_{2}}+t\|y\|_{E_{1}}\right)\text{.}$ (3)
For $x\in E_{2}$ and the parameters $0<\theta<1$ and $1\leq r<\infty$, this enables us to set
$\|x\|_{[E_{2},E_{1}]_{\theta,r}}:=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,E_{2},E_{1})\mathrm{d}t\right)^{\frac{1}{r}}$ (4)
and similarly, for $r=\infty$, we let
$\|x\|_{[E_{2},E_{1}]_{\theta,\infty}}:=\sup_{t>0}\left(t^{-\theta}K(x,t,E_{2},E_{1})\right)\text{.}$ (5)
Now the interpolation space of $E_{1}$ and $E_{2}$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined as
$[E_{2},E_{1}]_{\theta,r}:=\{x\in E_{2}\,:\,\|x\|_{[E_{2},E_{1}]_{\theta,r}}<\infty\}$ (6)
and this space is equipped with the $\|\,\cdot\,\|_{[E_{2},E_{1}]_{\theta,r}}$-norm. These spaces are again Banach spaces, see [3, Proposition 1.8 in Chapter 5]. In fact, such interpolation spaces perform a kind of interpolation of Banach spaces, as they are “in between” $E_{1}$ and $E_{2}$ in the sense that $E_{1}\hookrightarrow[E_{2},E_{1}]_{\theta,r}\hookrightarrow E_{1}$, see Lemma A.2.
Note that the definition of real interpolation spaces presented here is just one of several alternatives. Other ways, such as the $J$-method, yield equivalent norms. In certain situations, such equivalent norms can be helpful to gain insight into interpolation spaces.
Reproducing kernel Hilbert spaces and Banach spaces of functions
Let $X$ be a set and suppose that $E$ is a Banach space that consists of functions $f:X\rightarrow\mathbb{R}$. If for all $x\in X$, the evaluation maps
$\delta_{x}:E\rightarrow\mathbb{R},\quad f\mapsto f(x)$ (7)
are bounded functionals on $E$, then $E$ is called a Banach space of functions (BSF), see [19] for more details. These spaces recently attracted interest, as they yield insights into neural networks [2].
If additionally, the norm of such a Banach space of functions $H$ is induced by an inner product, for each $x\in X$, we can find a function $k(\cdot,x)\in H$ that is associated to $\delta_{x}$ by the Riesz representation theorem. Then $H$ is a reproducing kernel Hilbert space (RKHS) on $X$ with kernel $k$. We refer to [16, Chapter 4] for fundamental results on RKHS.
Now let $(X,\mathcal{A},\mu)$ be a measure space. Suppose that $H$ is an RKHS of a measurable and bounded kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$, that is the “embedding” $I_{k}:H\rightarrow\mathrm{L}^{2}(\nu)$ given by $f\mapsto[f]_{\sim}$ is a compact operator. In this case, we consider the self-adjoint, positive and compact operator $T_{k}:=I_{k}I_{k}^{*}$, called the integral operator of $k$, as it is given by
$T_{k}:\mathrm{L}^{2}(\nu)\rightarrow\mathrm{L}^{2}(\nu),\quad f\mapsto\left[\int_{X}k(\cdot,y)f(y)\mathrm{d}\nu(y)\right]_{\sim}\text{,}$ (8)
see e.g. [18, Lemma 2.2]. Let us denote the non-zero eigenvalues of $T_{k}$ in non-increasing order by $\mu=(\mu_{i})_{i\in I}$, where $I=\{1,...,n\}$ or $I=\mathbb{N}$. Due to the positivity and compactness of $T_{k}$, we know that $\mu_{i}>0$ and, in case that $I$ is infinite, $\lim_{i\rightarrow\infty}\mu_{i}=0$.
As proven in [18, Lemma 2.12], in such a situation, there are functions $e_{i}\in H$ such that $([e_{i}]_{\sim})_{i\in I}\subset\mathrm{L}^{2}(\nu)$ forms an ONS in $\mathrm{L}^{2}(\nu)$ consisting of eigenvectors of the integral operator $T_{k}$, such that additionally $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ forms an ONS inside $H$.
The range of $I_{k}$, that is $H]_{\sim}$, becomes a Hilbert space by using the quotient norm $\|[f]_{\sim}\|:=\inf\{\|g\|_{H}\,|\,I_{k}g=[f]_{\sim}\}$.
We summarise the above in the following assumption:
Assumption H. Suppose that $(X,\mathcal{A},\mu)$ is a measure space and that $H$ is an RKHS of a measurable kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$. Then, for a suitable index set $I=\{1,...,n\}$ or $I=\mathbb{N}$, we denote the eigenvalues of $T_{k}:\mathrm{L}^{2}(\nu)\rightarrow\mathrm{L}^{2}(\nu)$ by $\mu_{i}$ and the associated eigenfunctions by $e_{i}$ for $i\in I$.
Assumption H+. Let Assumption H be satisfied. In addition, let the measure space $(X,\mathcal{A},\mu)$ be complete and $\mu$ be $\sigma$-finite. Furthermore, assume that $k$ is a bounded kernel.
Power spaces
Our goal is to describe the interpolation spaces from (2) together with their norms by a spectral representation, i.e. in terms of eigenvalues and eigenfunctions of the operator $T_{k}$. To this end, for a “weight function” $w:I\rightarrow(0,\infty)$, we consider the following partition of $I$:
$M_{j}:=\left\{i\in I\,\middle|\,2^{j}<w(i)\leq 2^{j+1}\right\},\quad j\in\mathbb{Z}\text{.}$ (9)
Definition 2.1.
Let $(\mu_{i})_{i\in I}$ be a non-increasing sequence of positive numbers and $b=(b_{i})_{i\in I}$ be a sequence of real numbers. Then for $\theta>0$ and $1\leq r\leq\infty$, let
$c_{j}(b):=\left(\sum_{i\in M_{j}}b_{i}^{2}\mu_{i}^{-\theta}\right)^{\frac{1}{2}}$ (10)
and
$\|b\|_{(\mu,\theta,r)}:=\|(c_{j}(b))_{j\in\mathbb{Z}}\|_{\ell^{r}}\text{.}$ (11)
Furthermore, let $\ell^{(\mu,\theta,r)}:=\left\{b\,:\,\|b\|_{(\mu,\theta,r)}<\infty\right\}$ be equipped with the $\|\,\cdot\,\|_{(\mu,\theta,r)}$-norm.
It is not hard to see that, as the $\mu_{i}$ are positive, $\|\cdot\|_{(\mu,\theta,r)}$ is in fact a norm on $\ell^{(\mu,\theta,r)}$. The choice of powers of $2$ in (9) is for notational convenience. Replacing it by any $s>1$ yields equivalent norms. It will turn out that $\ell^{(\mu,\theta,r)}$ is actually a Banach space, see Remark 5.8. Note that in the case $r=2$, one finds that $\ell^{(\mu,\theta,r)}=\ell^{2}(\mu^{-\theta})$ with equal norms. In consequence, the following definition generalises the notion of a power of an RKHS from [18, Section 4].
Definition 2.2.
Let $H$, $e_{i}$ and $\mu_{i}$ be as introduced in Assumption H. Then the power space of $H$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined by
$H_{\nu}^{(\theta,r)}:=\left\{f\in\mathrm{L}^{2}(\nu)\,\middle|\,\exists b\in\ell^{(\mu,\theta,r)}\text{ with }f=\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\}$ (12)
with norm given by
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}:=\|b\|_{(\mu,\theta,r)}\text{.}$ (13)
Note that this norm is well defined due to the uniqueness of the Fourier coefficients $b_{i}$. Recall that in [18, Section 4] only the spaces $H_{\nu}^{\theta,2}$ were introduced and investigated. Moreover, for $r\neq 2$, the spaces $H_{\nu}^{(\theta,r)}$ are no longer Hilbert spaces.
3 Results
Our three main results are provided in the following. The first result states the norm equivalence of the power spaces $H_{\nu}^{\theta,r}$ to the interpolation spaces of the form (2), as previously mentioned. This is the foundation for the subsequent result, which shows that we can view these interpolation spaces as a BSF, in case they are embedded into $\mathrm{L}^{\infty}(\nu)$. The third result then analyses in which situations such an embedding exists.
Theorem 3.1.
Let $0<\theta<1$ and $1\leq r\leq\infty$ and suppose that $H$ is as in Assumption H. Then
$H_{\nu}^{(\theta,r)}=[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ (14)
and the norms on these spaces are equivalent.
Note that as $H]_{\sim}$ and $\mathrm{L}^{2}(\nu)$ consist of equivalence classes of functions, so does the interpolation space considered in Theorem 3.1. Not only for applications in learning theory, it is of interest whether such spaces can be viewed as actual function spaces with continuous evaluations functionals by a suitable choice of representatives. Note that for $r\neq 2$, the spaces $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ are no Hilbert spaces any more, so we have to expect BSFs instead of RKHSs.
Theorem 3.2.
Let $(X,\mathcal{A},\nu)$ and $H$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then there is a BSF $\overline{H}_{\nu}^{(\theta,r)}$ on $X$ such that the operator
$\overline{H}_{\nu}^{(\theta,r)}\rightarrow H_{\nu}^{(\theta,r)},\quad f\mapsto[f]_{\sim}$ (15)
is a well-defined isometric isomorphism with
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (16)
The existence of an embedding of the power space $H_{\nu}^{(\theta,r)}$ into $\mathrm{L}^{\infty}(\nu)$ seems to be a crucial requirement of Theorem 3.2. Let us elaborate in which situations we can expect such an embedding. For fine-index $r=2$, this question has been approached in [18, Section 5]. We extend this approach to $1\leq r\leq\infty$.
Theorem 3.3.
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$, where $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Then the following statements are equivalent:
  1. (i)
    It holds that
    $\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa$ (17)
  2. (ii)
    The embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous with operator norm at most $\kappa$.
4 Application to statistical learning
Let us introduce some basic notions from statistical learning theory, where we refer to [16] for details. We assume that $(X,\mathcal{A})$ is a measurable space, and that $Y\subset\mathbb{R}$ is bounded. Suppose there is a probability measure $P$ on $X\times Y$, wherefrom only some samples $(x_{1},y_{1}),...,(x_{n},y_{n})$ are known. Typically, one searches for a prediction function $f:X\rightarrow\mathbb{R}$ that predicts an $Y$-value associated to a given $X$-value with low risk:
$\mathcal{R}_{L,P}(f):=\int_{X\times Y}L(y,f(x))dP(x,y)\text{.}$ (18)
Hereby, $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is the loss function, such as the least squares loss $L(y,t)=(y-t)^{2}$. A function which minimises the risk over all prediction functions $f:X\rightarrow\mathbb{R}$ is called Bayes decision function and denoted by $f_{L,P}^{*}$.
To find a good prediction function, one can choose an RKHS $H$ on $X$, a $\lambda>0$ and perform a Tikhonov-regularised minimisation of the risk, that is
$f_{P,\lambda}\in\mathrm{argmin}_{f\in F}\;\lambda\|f\|_{H}^{2}+\mathcal{R}_{L,P}(f)$ (19)
where $P$ is in practice replaced by the empirical probability measure assigned to the given data. As done in [6], we assume in this section that the kernel of $H$ is bound by $1$.
Assuming that $L$ is the least squares loss, $f_{P,\lambda}$ can alternatively be found by the eigen-decomposition of $H$, see [18, Theorem 5.4]:
$f_{P,\lambda}=\sum_{i\in I}\frac{\mu_{i}}{\lambda+\mu_{i}}\left\langle f_{L,P}^{*}\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(P_{X})}e_{i}$ (20)
where we assume that $P_{X}$ is the marginal distribution, that Y=[1,1]Y=[-1,1] and that $H$, the eigenvalues $\mu_{i}$ and the basis $e_{i}$ are as in Assumption H. By this reformulation, we can to describe various other methods of regularised minimisation than (19) by replacing the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ by a more general function $\mu_{i}g_{\lambda}(\mu_{i})$. The following conditions taken from [6] ensure that a reasonable algorithm emerges.
Definition 4.1.
A family of functions $g_{\lambda}:[0,1]\rightarrow\mathbb{R}$ for $0<\lambda\leq 1$ satisfying
$\sup_{0<\mu\leq 1}\sup_{0\leq\lambda\leq 1}|\lambda g_{\lambda}(\mu)|$ $<\infty$ (21)
$\sup_{0<\mu\leq 1}\sup_{0\leq\lambda\leq 1}|\mu g_{\lambda}(\mu)|$ $<\infty$ (22)
$\lim_{\lambda\rightarrow 0}\mu g_{\lambda}(\mu)$ $=1$ (23)
for all $0\leq\mu\leq 1$ is called an admissible filter function, if there is a constant $\alpha_{0}>0$ such that for all $0<\alpha\leq\alpha_{0}$
$\sup_{0<\mu\leq 1}\sup_{0\leq\lambda\leq 1}|1-\mu g_{\lambda}(\mu)|\mu^{\alpha}\lambda^{-\alpha}<\infty\text{.}$ (24)
Let us now replace the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ in (20) by the more general $\mu_{i}g_{\lambda}(\mu_{i})$, that is
$f_{P,\lambda,g}:=\sum_{i\in I}\mu_{i}g_{\lambda}(\mu_{i})\left\langle f_{L,P}^{*}\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(P_{X})}e_{i}\text{.}$ (25)
This converges in $\mathrm{L}^{2}(P_{X})$ if $f_{L,P}^{*}\in\mathrm{L}^{2}(P_{X})$, as $\mu_{i}g_{\lambda}(\mu_{i})$ is bounded by (22). For example, to describe the Landweber iteration, that is a form of gradient descent with early stopping, we can choose
$g_{t}(\mu):=\sum_{i=0}^{t-1}(1-\mu)^{i}$ (26)
by using $\lambda:=t^{-1}$ with $t\in\mathbb{N}$. For details on the Landweber iteration and for further examples of various algorithms arising from different choices of $g_{\lambda}$, see [6].
We can estimate the “approximation error”, in our case this is $f_{P,\lambda,g}-f_{L,P}^{*}$ measured in interpolation space norms. This can be viewed as the distance of the optimal Bayes decision function $f_{L,P}^{*}$ to the result of the regularised empirical risk minimisation $f_{P,\lambda,g}$ in the case of full knowledge of the distribution $P$.
Proposition 4.2.
Let $X$ be a measurable space, $P$ a measure on $X\times[-1,1]$ and $k$ a measurable kernel on $X$ with RKHS $H$ that is compactly embedded into $\mathrm{L}^{2}(P_{X})$, where the eigenvalues of $T_{k}$ are bounded by $1$. Let $0<\beta<1$, $0<\theta<1$ and $1\leq r\leq\infty$. If $f_{L,P}^{*}\in[\mathrm{L}^{2}(P_{X}),[H]_{\sim}]_{\beta,r}$, the following holds.
  1. (i)
    There is a constant $C_{1}$ independent of $\lambda$ and $g_{\lambda}$ such that
    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{1}\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{.}$ (27)
  2. (ii)
    There is a constant $C_{2}$ independent of $\lambda$ and $g_{\lambda}$ such that
    $\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{2}\sup_{i\in I}\left(\big{|}\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{.}$ (28)
Suppose that in addition to the requirements of Proposition 4.2, we know that the embedding $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{\theta,r}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ is continuous with operator norm $\kappa$, that is
$\|[f_{P,\lambda,g}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ (29)
which yields in combination with (28) that
$\|[f_{P,\lambda,g}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}\leq C_{3}\sup_{i\in I}\left(\big{|}\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{.}$ (30)
Such estimates are play an important role in the statistical analysis of kernel-based learning algorithms, see e.g. [16, Theorem 7.23]. Note that the required existence of the embedding to $\mathrm{L}^{\infty}(\nu)$ may be checked via Theorem 3.3.
Remark 4.3.
We elaborate Proposition 4.2 for two concrete regularisation settings. To this end, we assume that $\beta>\theta$, and hence $\gamma:=\frac{\beta-\theta}{2}>0$.
  1. (i)
    Tikhonov regularisation. Here we have $g_{\lambda}(\mu)=\frac{1}{\lambda+\mu}$. For the supremum in Proposition 4.2, we find that
    $\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\gamma}\right)=\sup_{i\in I}\left(\frac{\lambda\mu_{i}^{\gamma}}{\lambda+\mu_{i}}\right)$ $\leq\sup_{0<\mu\leq 1}\left(\frac{\lambda\mu^{\gamma}}{\lambda+\mu}\right)$ (31)
    $\leq\lambda^{\gamma}\frac{\left(\frac{\gamma}{1-\gamma}\right)^{\gamma}}{1+\frac{\gamma}{1-\gamma}}$ (32)
    $=\lambda^{\gamma}\gamma^{\gamma}(1-\gamma)^{1-\gamma}$ (33)
    which can be seen by finding the zero of the derivative of the expression in question. Hence for the estimate in Proposition 4.2, we get
    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\lambda^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}$ (34)
    where we additionally use that $\gamma^{\gamma}(1-\gamma)^{1-\gamma}\leq 1$.
  2. (ii)
    Landweber iteration. We have $g_{t}(\mu)=\sum_{i=0}^{t-1}(1-\mu)^{i}$ with $\lambda=t^{-1}$. For $\mu<1$ we reformulate this by the geometric sum as $g_{t}=\frac{1-(1-\mu)^{t}}{\mu}$ to find that
    $\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\gamma}\right)=\sup_{i\in I}(1-\mu_{i})^{t}\mu_{i}^{\gamma}\leq\sup_{0<\mu<1}(1-\mu)^{t}\mu^{\gamma}=\left(\frac{t}{\gamma+t}\right)^{t}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}$ (35)
    which can be seen by finding the zero of the derivative of the expression in question. Hence for the estimate in Proposition 4.2, we get
    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{,}$ (36)
    where we note that estimating $\left(\frac{t}{\gamma+t}\right)^{t}\leq 1$ does not affect asymptotic properties in $t$.
5 Proofs
We can describe interpolation spaces of weighted $\mathrm{L}^{p}$-spaces by an equivalent norm via the following theorem. For the sake of completeness, we cite this result using our notation here.
Theorem 5.1.
Let $(X,\mathcal{A},\nu)$ be a measure space, and $w:X\rightarrow[\epsilon,\infty)$ be measurable, where $\epsilon>0$ is some constant. Furthermore, let $s>1$, $\theta\in(0,1)$, and $r\in[1,\infty)$. Then
$\|f\|=\left(\sum_{j\in\mathbb{Z}}\left(s^{-j\theta}\|f\cdot\mathbbm{1}_{\tilde{M}_{j}}\|_{\mathrm{L}^{p}(\nu)}\right)^{r}\right)^{\frac{1}{r}}\text{,}$ (37)
where $\tilde{M}_{j}:=\{x\in X\,|\,s^{-j}<w(x)^{\frac{1}{p}}\leq s^{-j+1}\}$, is an equivalent norm on $\mathrm{L}^{p}(\nu),\allowbreak\mathrm{L}^{p}(w)]_{(\theta,r)}$. An analogous result holds for $r=\infty$.
Proof.
The assertion directly follows from [7, Theorem 3.3], where we set $\mathscr{X}$ as $\mathrm{L}^{p}(\nu)$ and used the weight $\omega=w^{\frac{1}{p}}$, as the author applies the weight to the functions instead of the measures as we do here. Finally, we need to recall that the $K$-method and $J$-method used in [7] yield equivalent norms on the interpolation space. ∎
Let us adapt this result to our needs. To this end, we generalise the definition of $M_{j}$ from (9) to a general measure space $(X,\mathcal{A},\nu)$ and a measurable “weight” function $w:X\rightarrow(0,\infty)$:
$M_{j}:=\left\{x\in X\,\middle|\,\frac{1}{2^{j+1}}\leq\frac{1}{w(x)}<\frac{1}{2^{j}}\right\}\text{.}$ (38)
Lemma 5.2.
Let $(X,\mathcal{A},\nu)$ be a measure space and let $\varepsilon>0$, and $w:X\rightarrow[\epsilon,\infty)$ be measurable. Furthermore, let $\theta\in(0,1)$, $r\in[1,\infty)$, and let
$c_{j}(f):=\left(\int_{M_{j}}|f|^{p}w^{\theta}\mathrm{d}\nu\right)^{\frac{1}{p}}$ (39)
for $f\in[\mathrm{L}^{p}(\nu),\mathrm{L}^{p}(w)]_{\theta,r}$ and $j\in\mathbb{Z}$. Then an equivalent norm on $\mathrm{L}^{p}(\nu),\mathrm{L}^{p}(w)]_{\theta,r}$ is given by
$\|f\|:=\|(c_{j}(f))_{j\in\mathbb{Z}}\|_{\ell^{r}}\text{.}$ (40)
Proof.
[…]
$\tilde{M}_{-j}=\{x\in X\,|\,s^{j}<w(x)^{\frac{1}{p}}\leq s^{j+1}\}=\{x\in X\,|\,2^{-(j+1)}\leq w(x)^{-1}<2^{-j}\}=M_{j}\text{.}$ (41)
Hence performing an index swap from $j$ to j-j yields
[…]
if $1\leq r<\infty$, and similarly for $r=\infty$ using the supremum.
[…]
Auxiliary results on RKHS
For an RKHS $H$ as in Assumption H, we considered the space of equivalence classes $H]_{\sim}$ equipped with the quotient norm. It can be understood by the eigenfunction decomposition via the following lemma.
Lemma 5.3.
Let $H$, $\mu_{i}$ and $e_{i}$ be as in Assumption H. Then the mapping defined by
$\phi:[H]_{\sim}\rightarrow\ell^{2}(\mu^{-1}),\quad[f]_{\sim}\mapsto\left(\left\langle[f]_{\sim}\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(\nu)}\right)_{i\in I}$ (43)
is an isometric isomorphism.
Proof.
Recall that $(\sqrt{\mu_{i}}e_{i})_{i}$ is an ONS in $H$. We write $H_{I}:=\overline{\operatorname{\mathrm{span}}\{\sqrt{\mu_{i}}e_{i}|i\in I\}}$ for the subspace spanned by this ONS. By [18, Formula (17)], we know that $H_{I}=\overline{\text{Im}I_{k}^{*}}$, and hence [18, Formula (19)] shows $H_{I}^{\perp}=\text{Ker}I_{k}$. If we choose an ONB $(\tilde{e}_{j})_{j}$ of $H_{I}^{\perp}$ on a new index set $J$, we thus find
$[H]_{\sim}$ $=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}+\sum_{j\in J}b_{i}\tilde{e}_{j}\right]_{\sim}\,\middle|\,(a_{i})\in\ell^{2}(I),(b_{j})\in\ell^{2}(J)\right\}$ (44)
$=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]\,\middle|\,(a_{i})\in\ell^{2}(I)\right\}$ (45)
$=[H_{I}]_{\sim}\text{.}$ (46)
Since $\tilde{I}_{k}:=I_{k}|_{H_{I}\rightarrow[H]_{\sim}}$ is bijective by the considerations above, we see that for all $f\in[H_{I}]_{\sim}$ there is exactly one $h_{f}\in H_{I}$ such that $h_{f}]_{\sim}=\tilde{I}_{k}h_{f}=f$. This gives
$\|f\|_{[H_{I}]_{\sim}}=\inf\{\|h\|_{H}\,:\,h\in H_{I}\text{ with }[h]_{\sim}=f\}=\|h_{f}\|_{H}\text{.}$ (47)
Moreover, for $f\in[H]_{\sim}$ we find
$\|f\|_{[H]_{\sim}}^{2}$ $=\inf\{\|h\|_{H}^{2}\,:\,h\in H\text{ with }[h]_{\sim}=f\}$ (48)
$=\inf\{\|h_{I}+h_{0}\|_{H}^{2}\,:\,h_{I}\in H_{I}\text{ and }h_{0}\in\text{Ker}I_{k}\text{ with }[h_{I}+h_{0}]_{\sim}=f\}$ (49)
$=\inf\{\|h_{I}\|^{2}_{H}+\|h_{0}\|_{H}^{2}\,:\,h_{I}\in H_{I}\text{ and }h_{0}\in\text{Ker}I_{k}\text{ with }[h_{I}+h_{0}]_{\sim}=f\}$ (50)
$=\inf\{\|h_{I}\|^{2}_{H}\,:\,h_{I}\in H_{I}\text{ with }I_{k}h_{I}=f\}$ (51)
$=\|h_{f}\|^{2}_{H}\text{.}$ (52)
Combining both considerations, we find $\|\,\cdot\,\|_{[H]_{\sim}}=\|\,\cdot\,\|_{[H_{I}]_{\sim}}$.
Now for $h\in H_{I}$, let $a_{i}:=\left\langle h\,{,}\,\sqrt{\mu}_{i}e_{i}\right\rangle_{H}$ and $b_{i}:=\left\langle[h]_{\sim}\,{,}\,([e_{i}]_{\sim})_{i\in I}\right\rangle_{\mathrm{L}^{2}(\nu)}$. Then we have
$\sum_{i\in I}b_{i}[e_{i}]_{\sim}=[h]_{\sim}=\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]_{\sim}=\sum_{i\in I}a_{i}\sqrt{\mu_{i}}[e_{i}]_{\sim}\text{,}$ (53)
and thus $a_{i}\sqrt{\mu_{i}}=b_{i}$. Since $(a_{i})\in\ell^{2}(I)$, we get
$\|b\|_{\ell^{2}(\mu^{-1})}^{2}=\sum_{i\in I}b_{i}^{2}\mu_{i}^{-1}=\sum_{i\in I}\left(\frac{b_{i}}{\sqrt{\mu_{i}}}\right)^{2}=\sum_{i\in I}a_{i}^{2}=\|h\|_{H_{I}}^{2}\text{.}$ (54)
Consequently, the map $h\mapsto\left\langle[h]_{\sim}\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(\nu)}$ is isometric. By the arguments above, it is not hard to see that it is also surjective, thus it is an isometric isomorphism. ∎
The following lemma translates the interpolation space norm of $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ to the eigenfunction decomposition of the RKHS.
Lemma 5.4.
For $0<\theta<1$, $1\leq r\leq\infty$ and $\nu$, $H$, $e_{i}$ and $\mu_{i}$ as in Assumption H, the following map is an isometric isomorphism:
$[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}\rightarrow[\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)},\quad f\mapsto(\left\langle f\,{,}\,[e_{i}]_{\sim}\right\rangle)_{i\in I}$ (55)
Proof.
Let us define $\mathcal{H}:=\overline{[H]_{\sim}}^{\,\mathrm{L}^{2}(\nu)}$ and equip this space with the $\mathrm{L}^{2}(\nu)$-inner product. An application of Lemma A.3 to this orthogonal decomposition yields
$\left[\mathrm{L}^{2}(\nu),[H]_{\sim}\right]_{\theta,r}=\left[\mathcal{H}\oplus\mathcal{H}^{\perp},[H]_{\sim}\right]_{\theta,r}=\left[\mathcal{H},[H]_{\sim}\right]_{\theta,r}$ (56)
as $H]_{\sim}\hookrightarrow\mathcal{H}$. Now let us consider the operator
$\phi:\mathcal{H}\rightarrow\ell^{2}(I),\quad x\mapsto\left(\left\langle x\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(\nu)}\right)_{i\in I}\text{.}$ (57)
Note that this operator is an isometric isomorphism. The restricted operator
$\phi|_{[H]_{\sim}\rightarrow\ell^{2}(\mu^{-1})}$ (58)
is again an isometric isomorphism by Lemma 5.3. Hence Lemma A.1 shows that the operator restricted to the interpolation spaces
$\phi|_{\left[\mathcal{H},[H]_{\sim}\right]_{\theta,r}\rightarrow\left[\ell^{2}(I),\ell^{2}(\mu^{-1})\right]}$ (59)
is again an isometric isomorphism. This proves the assertion. ∎
Proof of Theorem 3.1
Proof of Theorem 3.1.
We apply Lemma 5.2 to the counting measure on $I$ and the weight sequence $\mu^{-1}$ to find that the identity mapping $J:[\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}\rightarrow\ell^{(\mu,\theta,r)}$ is a norm equivalence.
Consider the isometric isomorphism $\alpha:[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{\theta,r}\rightarrow[\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}$ from Lemma 5.4. It maps functions to their $\mathrm{L}^{2}(\nu)$-Fourier coefficients with respect to the ONS $e_{i}$.
The map $\beta:\ell^{(\mu,\theta,r)}\rightarrow H_{\nu}^{(\theta,r)}$ given by $b\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ is an isometric isomorphism by Definition 2.2.
We note that for $f\in[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{\theta,r}$, we have
$\beta(J(\alpha(f)))=\beta(\alpha(f))=\beta((\left\langle f\,{,}\,[e_{i}]_{\sim}\right\rangle)_{i\in I})=\sum_{i\in I}\left\langle f\,{,}\,[e_{i}]_{\sim}\right\rangle[e_{i}]_{\sim}=f\text{,}$ (60)
which shows that the following diagram commutes:
[…]
This proves the assertion. ∎
Remark 5.5.
By Lemma A.2, we know that $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$. Hence we can apply Theorem 3.1 to see that for the power spaces, we have $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$ as well. However, the operator norms of these two embedding operators might differ by a constant, as the norms are equivalent, but not necessarily equal.
Lemma 5.6.
Let $0<\theta<1$ and $1\leq r\leq\infty$. If $b\in\ell^{(\mu,\theta,r)}$, then $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in the power space $H_{\nu}^{(\theta,r)}$.
Proof.
For finite $I$, the assertion is clear. Hence let us assume $I=\mathbb{N}$. Let $f:=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$, where convergence is in $\mathrm{L}^{2}(\nu)$ at first. To show convergence in $H_{\nu}^{(\theta,r)}$-norm, we notice that $\ell^{(\mu,\theta,r)}\triangleq[\ell^{2}(I),\allowbreak\ell^{2}(\mu^{-1})]_{(\theta,r)}$ by Lemma 5.2. Applying Lemma A.2 yields $\ell^{(\mu,\theta,r)}\hookrightarrow\ell^{2}(I)$. If $C$ denotes the corresponding operator norm, we thus find
$\left\|f-\sum_{i=1}^{n}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{(\mu,\theta,r)}\leq C\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{\ell^{2}(I)}\xrightarrow{n\rightarrow\infty}0\text{.}$ (61)
Liftings
Let us introduce a measure-theoretic technique that will play a role in the subsequent proofs. To this end, we consider the map
$\mathcal{L}^{\infty}(X)\rightarrow\mathrm{L}^{\infty}(\nu),\quad f\mapsto[f]_{\sim}$ (62)
which assigns to each $f$ the equivalence class of functions. This mapping is surjective by definition of $\mathrm{L}^{\infty}(\nu)$, hence it possesses a, not necessarily unique, right inverse. However, this reasoning does not yield further properties of such a right inverse. Nonetheless, by means of lifting theory, one can show that a bounded linear right inverse exists.
Theorem 5.7.
Let $\nu$ be a $\sigma$-finite measure on the measurable space $(X,\mathcal{A})$ such that $\mathcal{A}$ is $\nu$-complete. Then there exists a bounded, linear, and injective operator
$\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ (63)
that satisfies $\|\phi\|\leq 1$ and
$[\phi([f]_{\sim})]_{\sim}=[f]_{\sim}$ (64)
for all $f\in\mathcal{L}^{\infty}(\nu)$.
The proof is based on lifting theory as developed in [8]. For a derivation of the specific result, see [18, Equation (58)].
So in other words, for each equivalence class in $\mathrm{L}^{\infty}(\nu)$, the operator $\phi$ chooses a bounded representative in a linear and continuous way.
Auxiliary results for the proofs of Theorems 3.2 and 3.3
Let us now introduce a setup which makes it possible to apply our results for direct sums of Banach spaces from Definition B.1.
Remark 5.8.
Let $0<\theta<1$ and $1\leq r\leq\infty$ and suppose that $\mu=(\mu_{i})_{i\in\mathbb{N}}$ is a non-increasing sequence of positive real numbers. Let us define the spaces $E_{j}:=\ell^{2}((\mu_{i}^{-\theta})_{i\in M_{j}})$.
Then for $j\in\mathbb{Z}$, let us consider the projection maps $\pi_{j}$, that map any sequence $(b_{i})_{i\in I}$ to the subsequence $(b_{i})_{i\in M_{j}}$. More specifically, we define
$\pi_{j}^{(r)}:\ell^{(\mu,\theta,r)}\rightarrow E_{j},\quad b\mapsto(b_{i})_{i\in M_{j}}\text{.}$ (65)
Then by Definition 2.1 and Appendix B, we know that
$\pi^{(r)}:\ell^{(\mu,\theta,r)}\rightarrow{\bigoplus_{j\in\mathbb{Z}}}^{(r)}E_{j},\quad b\mapsto(\pi_{j}^{(r)}(b))_{j\in\mathbb{Z}}$ (66)
is an isometric isomorphism, and hence the spaces $\ell^{(\mu,\theta,r)}$ are Banach spaces.
Now note that by the definition of the $E_{j}$-inner product, we have
$\left\langle\pi_{j}(b)\,{,}\,\pi_{j}(a)\right\rangle_{E_{j}}=\sum_{i\in M_{j}}a_{i}b_{i}\mu_{i}^{-\theta}$ (67)
for sequences $a,b$ of real numbers. Hence we find that
$\sum_{j\in\mathbb{Z}}\left\langle\pi_{j}(a)\,{,}\,\pi_{j}(b)\right\rangle_{E_{j}}=\sum_{j\in\mathbb{Z}}\sum_{i\in M_{j}}a_{i}b_{i}\mu_{i}^{-\theta}=\sum_{i\in I}a_{i}b_{i}\mu_{i}^{-\theta}$ (68)
whenever one of the sums converges absolutely. By Lemma B.3 and the isometric isomorphism (66), this can be viewed as a dual pairing of $\ell^{(\mu,\theta,r)}$ and $\ell^{(\mu,\theta,r^{\prime})}$ if $\frac{1}{r}+\frac{1}{r^{\prime}}=1$.
This setup now enables us to prove what can be viewed as a kind of Hölder inequality for the $\ell^{(\mu,\theta,r)}$-spaces.
Lemma 5.9.
Let $0<\theta<1$ and suppose that $1\leq r\leq\infty$ with $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Let $\mu=(\mu_{i})_{i\in I}$ be a non-increasing sequence of positive real numbers. Then for $b\in\ell^{(\mu,\theta,r)}$ and $a\in\ell^{(\mu,\theta,r^{\prime})}$ we have
$\sum_{i\in I}|a_{i}b_{i}|\mu_{i}^{-\theta}\leq\|a\|_{\ell^{(\mu,\theta,r^{\prime})}}\|b\|_{\ell^{(\mu,\theta,r)}}\text{.}$ (69)
Proof.
By Remark 5.8, for all $a\in\ell^{(\mu,\theta,r^{\prime})}$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
$\sum_{i\in I}|a_{i}b_{i}|\mu_{i}^{-\theta}$ $=\sum_{j\in\mathbb{Z}}\sum_{i\in M_{j}}|a_{i}b_{i}|\mu_{i}^{-\theta}$ (70)
$\leq\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}a_{i}^{2}\mu_{i}^{-\theta}\right)^{\frac{1}{2}}\left(\sum_{i\in M_{j}}b_{i}^{2}\mu_{i}^{-\theta}\right)^{\frac{1}{2}}$ (71)
$\leq\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}a_{i}^{2}\mu_{i}^{-\theta}\right)^{\frac{r^{\prime}}{2}}\right)^{\frac{1}{r^{\prime}}}\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}b_{i}^{2}\mu_{i}^{-\theta}\right)^{\frac{r}{2}}\right)^{\frac{1}{r}}$ (72)
$=\|a\|_{\ell^{(\mu,\theta,r^{\prime})}}\|b\|_{\ell^{(\mu,\theta,r)}}\text{,}$ (73)
where we first applied Cauchy-Schwarz and then Hölder’s inequality. ∎
The following lemma is essentially given by an application of Lemma B.4 to the setting introduced in Remark 5.8.
Lemma 5.10.
Let $0<\theta<1$ and suppose that $1\leq r\leq\infty$ with $\frac{1}{r}+\frac{1}{r^{\prime}}=1$. Let furthermore $\mu=(\mu_{i})_{i\in I}$ be a non-increasing sequence of positive real numbers and $a=(a_{i})_{i\in I}$ be a sequence of real numbers. Then the following statements are equivalent for all $C\geq 0$:
  1. (i)
    For all sequences $b\in\ell^{(\mu,\theta,r)}$, the sum $\sum_{i\in I}a_{i}b_{i}\mu_{i}^{-\theta}$ converges and it holds that
    $\left|\sum_{i\in I}a_{i}b_{i}\mu_{i}^{-\theta}\right|\leq C\|b\|_{\mu,\theta,r}\text{.}$ (74)
  2. (ii)
    It holds that $a\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|a\|_{(\mu,\theta,r^{\prime})}\leq C$.
Proof.
Let us use the notation from Remark 5.8. By (68), (i) means that for all $b\in\ell^{(\mu,\theta,r)}$, we have
$\left|\sum_{j\in\mathbb{Z}}\left\langle\pi_{j}(a)\,{,}\,\pi_{j}^{(r)}(b)\right\rangle_{E_{j}}\right|=\left|\sum_{i\in I}a_{i}b_{i}\mu_{i}^{-\theta}\right|\leq C\|b\|_{(\mu,\theta,r)}=C\|\pi^{(r)}(b)\|_{\bigoplus_{j\in\mathbb{Z}}^{(r)}E_{j}}$ (75)
with absolute convergence of the sums by Remark B.2. As $\pi^{(r)}:\ell^{(\mu,\theta,r)}\rightarrow{\bigoplus_{j\in\mathbb{Z}}}^{(r)}E_{j}$ is an isometric isomorphism, this means that for all $x\in{\bigoplus_{j\in\mathbb{Z}}}^{(r)}E_{j}$, it holds that
$\left|\sum_{j\in\mathbb{Z}}\left\langle\pi_{j}(a)\,{,}\,x_{j}\right\rangle_{E_{j}}\right|\leq C\|x\|_{\bigoplus_{j\in\mathbb{Z}}^{(r)}E_{j}}$ (76)
where we replaced $\pi^{(r)}(b)$ by $x$ and applied the isometry property on the right hand side. An application of Proposition B.5 shows that this is the case if and only if $(\pi_{j}(a))_{j\in\mathbb{Z}}\in\bigoplus_{j\in\mathbb{Z}}^{(r^{\prime})}E_{j}$ with norm $\|(\pi_{j}(a))_{j\in\mathbb{Z}}\|_{r^{\prime}}\leq C$. An application of the isometric isomorphism $\pi^{(r^{\prime})}$ finally shows that this is equivalent to $a\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|a\|_{(\mu,\theta,r^{\prime})}\leq C$, which is (ii). ∎
The following lemma gives some details on the construction of $\overline{H}_{\nu}^{(\theta,r)}$ in Theorem 3.2.
Lemma 5.11.
Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Let
$\overline{e}_{i}:=\phi([e_{i}]_{\sim})\text{,}$ (77)
where $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ is the lifting according to Theorem 5.7. Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous. Then the following statements are true.
  1. (i)
    The series $\sum_{i\in I}b_{i}\overline{e}_{i}$ converges uniformly for all $b\in\ell^{(\mu,\theta,r)}$.
  2. (ii)
    It holds that $\sum_{i\in I}b_{i}\overline{e}_{i}=\sum_{i\in I}b_{i}e_{i}$ almost everywhere for all $b\in\ell^{(\mu,\theta,r)}$.
  3. (iii)
    The space
    $\overline{H}_{\nu}^{(\theta,r)}:=\left\{\sum_{i\in I}b_{i}\overline{e}_{i}\,\middle|\,b\in\ell^{(\mu,\theta,r)}\right\}$ (78)
    is a BSF, where the BSF-norm is inherited from the $\ell^{(\mu,\theta,r)}$-norm of the Fourier coefficients $b_{i}$.
  4. (iv)
    The mapping
    $H_{\nu}^{(\theta,r)}\rightarrow\overline{H}_{\nu}^{(\theta,r)},\quad\sum_{i\in I}b_{i}[e_{i}]_{\sim}\mapsto\sum_{i\in I}b_{i}\overline{e}_{i}$ (79)
    is an isometric isomorphism.
  5. (v)
    It holds that
    $\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (80)
Proof.
As the RKHS $H$ has a bounded kernel, the eigenfunctions $e_{i}$ are bounded as well. Hence by Theorem 5.7, the lifted eigenfunctions $\overline{e}_{i}=\phi([e_{i}]_{\sim})$ are measurable, bounded and we know that for each $i\in I$, there is a $\nu$-null set $N_{i}$ such that $e_{i}(x)=\overline{e}_{i}(x)$ for all $x\in X\setminus N_{i}$. As $I$ is at most countable, $N:=\bigcup_{i\in I}N_{i}$ is again measurable with $\nu(N)=0$. By that, for all $x\in X\setminus N$ and all $i\in I$, we have $\overline{e}_{i}(x)=e_{i}(x)$. Furthermore, let us denote
$\kappa:=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (81)
(i). Let $b\in\ell^{(\mu,\theta,r)}$. Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H_{\nu}^{(\theta,r)}$. Due to the continuity of the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it converges in $\mathrm{L}^{\infty}(\nu)$ as well and we have
$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (82)
By continuity of the lifting operator $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ from Theorem 5.7, we thus find
$\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)=\sum_{i\in I}b_{i}\phi([e_{i}]_{\sim})=\sum_{i\in I}b_{i}\overline{e}_{i}\text{,}$ (83)
so $\sum_{i\in I}b_{i}\overline{e}_{i}$ converges in $\mathcal{L}^{\infty}(X)$, that is uniformly.
(ii). We know that $\overline{e}_{i}(x)=e_{i}(x)$ holds for all $i\in I$ and all $x\in X\setminus N$, which yields the assumption by (i).
(iii). As $\ell^{(\mu,\theta,r)}$ is a Banach space by Remark 5.8, it follows that $\overline{H}_{\nu}^{(\theta,r)}$ is a Banach space as well. By the uniform convergence in (i), $\overline{H}_{\nu}^{(\theta,r)}$ consists of functions. It remains to show that the evaluation functionals are continuous. To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$. Because $\phi$ has operator norm at most $1$, by (83) we find that
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)\right\|_{\mathcal{L}^{\infty}(X)}\leq\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\text{.}$ (84)
Now for all $x\in X$, we consider $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ to find that
$\left|\delta_{x}\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)\right|=\left|\sum_{i\in I}b_{i}\overline{e}_{i}(x)\right|\leq\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}$ (85)
where for second to last inequality, we combined (82) and (84).
(iv). This follows as the norms on $H_{\nu}^{(\theta,r)}$ and on $\overline{H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
(v). Note that by using (85), we find that
$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}=\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{H_{\nu}^{(\theta,r)}}\text{,}$ (86)
which shows $\leq$. For $\geq$, we note that
$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|f\|_{\mathcal{L}^{\infty}(X)}\geq\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|$ (87)
where the last step follows as $\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}=\|[f]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$. ∎
Proof of Theorem 3.2.
The assertion follows directly from Lemma 5.11 by using the inverse of the isometric isomorphism from (iv). ∎
Proof of Theorem 3.3.
First of all, note that by Definition 2.2, an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ can be written as a series
$[g]_{\sim}=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ (88)
for a suitable $b\in\ell^{(\mu,\theta,r)}$, where the convergence is in $H_{\nu}^{(\theta,r)}$ by Lemma 5.6. In addition, it holds that
$\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}=\|b\|_{(\mu,\theta,r)}\text{.}$ (89)
For the rest of this proof, we write $f_{i}(x):=e_{i}(x)\mu_{i}^{\theta}$ for all $i\in I$ and all $x\in X$.
“(i) $\Rightarrow$ (ii)”: By Property (i), there is a $\nu$-null set $N$ such that
$\left\|\left(f_{i}(x)\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa$ (90)
holds for all $x\in X\setminus N$. Hence it follows for $x\in X\setminus N$ and all $b\in\ell^{(\mu,\theta,r)}$ that
$\sum_{i\in I}|b_{i}e_{i}(x)|$ $=\sum_{i\in I}|b_{i}f_{i}(x)|\mu_{i}^{-\theta}$ (91)
$\leq\|(f_{i}(x))_{i\in I}\|_{(\mu,\theta,r^{\prime})}\|b\|_{(\mu,\theta,r)}$ (92)
$\leq\kappa\,\|b\|_{(\mu,\theta,r)}\text{,}$ (93)
[…]
Consequently, we find for such an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ that
$\|[g]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$ $=\inf\left\{B\geq 0\,:\,\exists M\text{ with }\nu(M)=0\text{ and }\sup_{x\in X\setminus M}\left|g(x)\right|\leq B\right\}$ (94)
$\leq\sup_{x\in X\setminus(N\cup\tilde{N})}\left|g(x)\right|$ (95)
$=\sup_{x\in X\setminus(N\cup\tilde{N})}\left|\sum_{i\in I}b_{i}e_{i}(x)\right|$ (96)
$\leq\kappa\|b\|_{(\mu,\theta,r)}$ (97)
$=\kappa\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$ (98)
by an application of (93). This proves (ii).
“(ii) $\Rightarrow$ (i)”: By Lemma 5.11, for almost all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
$\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}=\sum_{i\in I}b_{i}e_{i}(x)=\sum_{i\in I}b_{i}\overline{e}_{i}(x)\leq\kappa\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\overline{H}_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (99)
Thus, for almost all $x\in X$, an application of Lemma 5.10 to $a_{i}=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$. This finally shows (i). ∎
Proofs concerning applications
Lemma 5.12.
Suppose that $(\mu_{i})_{i\in I}$ is a non-increasing sequence of positive numbers, $0<\beta<1$, $0<\theta<1$ and $1\leq r\leq\infty$. Let $(\alpha_{i})_{i\in I}$ and $(b_{i})_{i\in I}$ be sequences of real numbers. Then it holds that
$\|(\alpha_{i}b_{i})_{i\in I}\|_{(\mu,\theta,r)}\leq\sup_{i\in I}\left(|\alpha_{i}|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}\text{.}$ (100)
Proof.
$\|(\alpha_{i}b_{i})_{i\in I}\|_{(\mu,\theta,r)}$ $=\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\big{(}\alpha_{i}b_{i}\big{)}^{2}\mu_{i}^{-\theta}\right)^{\frac{r}{2}}\right)^{\frac{1}{r}}$ (101)
$=\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\alpha_{i}^{2}\mu_{i}^{\beta-\theta}b_{i}^{2}\mu_{i}^{-\beta}\right)^{\frac{r}{2}}\right)^{\frac{1}{r}}$ (102)
$\leq\sup_{i\in I}\left(|\alpha_{i}|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}b_{i}^{2}\mu_{i}^{-\beta}\right)^{\frac{r}{2}}\right)^{\frac{1}{r}}$ (103)
$=\sup_{i\in I}\left(|\alpha_{i}|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (104)
Proof of Proposition 4.2.
For $i\in I$, we write $b_{i}:=\langle f_{L,P}^{*},[e_{i}]_{\sim}\rangle_{\mathrm{L}^{2}(P_{X})}$. Then we find that
$f_{L,P}^{*}=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ (105)
with convergence in $\mathrm{L}^{2}(P_{X})$ as $f^{*}_{L,P}\in\mathrm{L}^{2}(P_{X})$ by Remark 5.5. In addition, Theorem 3.1 yields a $C_{4}\geq 0$ independent of $f_{L,P}^{*}$ such that
$\|b\|_{(\mu,\beta,r)}\leq C_{4}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{.}$ (106)
Now note that we have $(\mu_{i}g_{\lambda}(\mu_{i})b_{i})_{i\in I}\in\ell^{2}(\mu^{-1})$ because $(b_{i})_{i\in I}$ is an $\ell^{2}$-sequence and $g_{\lambda}$ is bounded by Definition 4.1. Hence by (25), we see that
$f_{P,\lambda,g}=\sum_{i\in I}{\mu_{i}g_{\lambda}(\mu_{i})b_{i}}e_{i}$ (107)
converges in $H$. Hence it follows that
$[f_{P,\lambda,g}]_{\sim}=I_{k}f_{P,\lambda,g}=I_{k}\left(\sum_{i\in I}\mu_{i}g_{\lambda}(\mu_{i})b_{i}e_{i}\right)=\sum_{i\in I}\mu_{i}g_{\lambda}(\mu_{i})b_{i}[e_{i}]_{\sim}$ (108)
with convergence in $\mathrm{L}^{2}(P_{X})$.
(i). By the $\mathrm{L}^{2}(P_{X})$-orthogonal series of $f_{L,P}^{*}$ in (105) and the one of $f_{P,\lambda,g}]_{\sim}$ in (108), we conclude
$[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}=\sum_{i\in I}(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}[e_{i}]_{\sim}$ (109)
with convergence in $\mathrm{L}^{2}(P_{X})$.
Then by the norm equivalence from Theorem 3.1, we find a $C_{6}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
$\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{6}\|(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}\|_{(\mu,\theta,r)}$ (110)
$\leq C_{6}\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (111)
From this, the assertion follows by the upper bound on $\|b\|_{(\mu,\beta,r)}$ given in (106).
(ii). By the norm equivalence in Theorem 3.1, we find a $C_{5}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
$\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{5}\|\mu_{i}g_{\lambda}(\mu_{i})b_{i}\|_{(\mu,\theta,r)}$ (112)
$\leq C_{5}\sup_{i\in I}\left(|\mu_{i}g_{\lambda}(\mu_{i})|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (113)
From this, the estimate follows by the upper bound on $\|b\|_{(\mu,\beta,r)}$ given in (106). ∎
A Appendix: Interpolation spaces
In this appendix, we collect some auxiliary results on interpolation spaces, which are probably known to experts.
We begin by proving that isometric isomorphisms can be carried over from Banach spaces $E_{2}\hookrightarrow E_{1}$ to the interpolation spaces $E_{1},E_{2}]_{\theta,r}$. For a map $f:A\rightarrow B$ and $A^{\prime}\subset A$, $B^{\prime}\subset B$, we use the notation $f|_{A^{\prime}\rightarrow B^{\prime}}$ for the map $A^{\prime}\rightarrow B^{\prime}$ given by $x\mapsto f(x)$ in case it is well defined.
Lemma A.1.
Let $E_{2}\hookrightarrow E_{1}$ and $F_{2}\hookrightarrow F_{1}$ be Banach spaces. Suppose there is a bounded linear operator $\phi:E_{1}\rightarrow F_{1}$ such that $\phi|_{E_{1}\rightarrow F_{1}}$ and $\phi|_{E_{2}\rightarrow F_{2}}$ are isometric isomorphisms. Then for $0<\theta<1$ and $1\leq r\leq\infty$, the following map is an isometric isomorphism:
$\phi|_{[E_{1},E_{2}]_{\theta,r}\rightarrow[F_{1},F_{2}]_{\theta,r}}$ (114)
Proof.
Let us begin by showing that $\phi|_{[E_{1},E_{2}]_{\theta,r}\rightarrow[F_{1},F_{2}]_{\theta,r}}$ is an isometry. To this end note that for all $x\in E_{1}$ and $t>0$ we have
$K(x,t,E_{1},E_{2})$ $=\inf_{y\in E_{2}}\left(\|x-y\|_{E_{1}}+t\|y\|_{E_{2}}\right)$ (115)
$=\inf_{y\in E_{2}}\left(\|\phi(x-y)\|_{F_{1}}+t\|\phi(y)\|_{F_{2}}\right)$ (116)
$=\inf_{z\in F_{2}}\left(\|\phi(x)-z\|_{F_{1}}+t\|z\|_{F_{2}}\right)$ (117)
$=K(\phi(x),t,F_{1},F_{2})\text{.}$ (118)
This shows that $\|\phi(x)\|_{[F_{1},F_{2}]_{\theta,r}}=\|x\|_{[E_{1},E_{2}]_{\theta,r}}$, that is $\phi|_{[E_{1},E_{2}]_{\theta,r}\rightarrow[F_{1},F_{2}]_{\theta,r}}$ is isometric. Finally, applying the same argument to $\phi^{-1}_{E_{i}\rightarrow F_{i}}$ shows that $\phi^{-1}|_{[F_{1},F_{2}]_{\theta,r}\rightarrow[E_{1},E_{2}]_{\theta,r}}$ is also isometric. Obviously, it is the inverse to $\phi|_{[E_{1},E_{2}]_{\theta,r}\rightarrow[F_{1},F_{2}]_{\theta,r}}$. ∎
Let us collect some properties of interpolation spaces for the situation considered here, that is $E_{2}\hookrightarrow E_{1}$:
Lemma A.2.
Let $E_{2}\hookrightarrow E_{1}$ be Banach spaces. Suppose that $0<\theta_{1}\leq\theta_{2}<1$ and $1\leq r_{1}\leq r_{2}\leq\infty$. Then the following properties hold:
  1. (i)
    $E_{1},E_{2}]_{\theta_{1},r_{1}}\hookrightarrow E_{1}$
  2. (ii)
    $E_{2}\hookrightarrow[E_{1},E_{2}]_{\theta_{1},r_{1}}$
  3. (iii)
    $E_{1},E_{2}]_{\theta_{2},r_{1}}\hookrightarrow[E_{1},E_{2}]_{\theta_{1},r_{1}}$
  4. (iv)
    $E_{1},E_{2}]_{\theta_{1},r_{1}}\hookrightarrow[E_{1},E_{2}]_{\theta_{1},r_{2}}$
Proof.
(i). This follows by [3, Proposition 1.8 in Chapter 5] because we have $E_{1}+E_{2}=E_{1}$ due to the embedding $E_{2}\hookrightarrow E_{1}$.
(ii). This follows by [3, Proposition 1.8 in Chapter 5] because we have $E_{1}\cap E_{2}=E_{2}$ due to the embedding $E_{2}\hookrightarrow E_{1}$.
(iii). See [4, 3.4.1 (d)], where the notation is clarified in [4, page VII].
(iv). This follows by [3, Proposition 1.10 in Chapter 5]. ∎
Lemma A.3.
Let $H$ be a Hilbert space and let $H_{1}\subset H$ be a closed subspace. Suppose the Hilbert space $H_{2}$ satisfies $H_{2}\hookrightarrow H_{1}$. Then for $0<\theta<1$ and $1\leq r\leq\infty$, it holds that
$[H,H_{2}]_{\theta,r}=[H_{1},H_{2}]_{\theta,r}$ (119)
and the norms of both spaces are equal.
Proof.
Our first goal is to show
$[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}\subset H_{1}$ (120)
To this end, we first note that $H_{1},H_{2}]_{\theta,r}\subset H_{1}$ by Lemma A.2.
For the proof of $H,H_{2}]_{\theta,r}\subset H_{1}$ we pick an $x\in[H,H_{2}]_{\theta,r}$. As the latter space is included in $H=H_{1}\oplus H_{1}^{\perp}$ by Lemma A.2, we consider the unique decomposition $x=x_{1}+x_{0}$ with $x_{1}\in H_{1}$ and $x_{0}\in H_{1}^{\perp}$. It suffices to show that $x_{0}=0$. To this end, we assume the converse, that is $\|x_{0}\|_{H}>0$. Then for $t>0$, it follows that
$K(x,t,H,H_{2})$ $=\inf_{y\in H_{2}}\left(\|x-y\|_{H}+t\|y\|_{H_{2}}\right)$ (121)
$=\inf_{y\in H_{2}}\left(\|x_{1}-y+x_{0}\|_{H}+t\|y\|_{H_{2}}\right)$ (122)
$=\inf_{y\in H_{2}}\left(\sqrt{\|x_{1}-y\|_{H}^{2}+\|x_{0}\|_{H}^{2}}+t\|y\|_{H_{2}}\right)$ (123)
$\geq\|x_{0}\|_{H}\text{.}$ (124)
Hence for $1\leq r<\infty$, we find that $x\notin[H,H_{2}]_{\theta,r}$ as
$\|x\|_{[H,H_{2}]_{\theta,r}}$ $=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,H,H_{2})\mathrm{d}t\right)^{\frac{1}{r}}$ (125)
$\geq\left(\int_{0}^{\infty}t^{-\theta r-1}\|x_{0}\|_{H}^{r}\mathrm{d}t\right)^{\frac{1}{r}}$ (126)
$=\infty\text{.}$ (127)
For $r=\infty$, we find the same result by estimating
$\|x\|_{[H,H_{2}]_{\theta,\infty}}=\sup_{t>0}\left(t^{-\theta}K(x,t,H,H_{2})\right)\geq\sup_{t>0}\left(t^{-\theta}\|x_{0}\|_{H}\right)=\infty\text{.}$ (128)
Hence we find a contradiction to $x\in[H,H_{2}]_{\theta,r}$. This shows the desired $x_{0}=0$.
Let us now fix an $x\in[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}$. By (120), we note that $x\in H_{1}$ and hence we find
$K(x,t,H,H_{2})$ $=\inf_{y\in H_{2}}\left(\|x-y\|_{H}+t\|y\|_{H_{2}}\right)$ (129)
$=\inf_{y\in H_{2}}\left(\|x-y\|_{H_{1}}+t\|y\|_{H_{2}}\right)=K(x,t,H_{1},H_{2})\text{.}$ (130)
The equality of the interpolation spaces and their norms hence follows. ∎
B Appendix: Direct sums of Banach spaces
A closer look at Definition 2.1 reveals that the $\ell^{(\mu,\theta,r)}$-norms are built up from two nested norms: The inner part (10) of the $(\mu,\theta,r)$-norm originates from an $\ell^{2}(\mu^{-\theta})$-norm taken over the restricted indices in $M_{j}$, while the outer sum in (11) corresponds to the $\ell^{r}$-norm taken over the values given by the inner part.
To deal with these norms, let us introduce such nested norms in a more general setting. For notational convenience, let us assume that $I=\{1,...,n\}$ or $I=\mathbb{N}$ is an index set.
Definition B.1.
Let $I$ be an index set and suppose that $E_{i}$ are Banach spaces for all $i\in I$. Then for $1\leq p\leq\infty$ we define the $p$-direct sum of the $E_{i}$ as
${\bigoplus_{i\in I}}^{(p)}E_{i}:=\left\{x=(x_{i})_{i\in I}\,\middle|\,x_{i}\in E_{i}\text{ and }\|x\|_{p}<\infty\right\}\text{,}$ (131)
where the norm is defined by $\|x\|_{p}:=\|(\|x_{i}\|_{E_{i}})_{i\in I}\|_{\ell^{p}}$. We define the 0-direct sum as
${\bigoplus_{i\in I}}^{(0)}E_{i}:=\left\{x=(x_{i})_{i\in I}\,\middle|\,x_{i}\in E_{i}\text{ and }\lim_{i\rightarrow\infty}\|x_{i}\|_{E_{i}}=0\right\}$ (132)
where the limit of a finite sequence is defined as 0. We equip this space with the norm $\|x\|_{0}:=\|x\|_{\infty}$.
It is well known that such $p$-direct sums of Banach spaces $E_{i}$ are again Banach spaces, see e.g. [1, Section 16.11].
Remark B.2.
Let $E_{i}$ be Banach spaces for indexes $i\in I$, let $1\leq p\leq\infty$, $C>0$ and suppose that $x_{i}^{\prime}\in E_{i}^{\prime}$ for all $i\in I$. Then the following holds:
$\sum_{i\in I}x_{i}^{\prime}(x_{i})\text{ converges for all }x\in{\bigoplus_{i\in I}}^{(p)}E_{i}\;$ $\Leftrightarrow\;\sum_{i\in I}|x_{i}^{\prime}(x_{i})|\text{ converges for all }x\in{\bigoplus_{i\in I}}^{(p)}E_{i}$ (133)
$\sum_{i\in I}x_{i}^{\prime}(x_{i})\leq C\|x\|_{p}\text{ for all }x\in{\bigoplus_{i\in I}}^{(p)}E_{i}\;$ $\Leftrightarrow\;\sum_{i\in I}|x_{i}^{\prime}(x_{i})|\leq C\|x\|_{p}\text{ for all }x\in{\bigoplus_{i\in I}}^{(p)}E_{i}$ (134)
$\sup_{\|x\|_{p}\leq 1}\sum_{i\in I}x_{i}^{\prime}(x_{i})\;$ $=\;\sup_{\|x\|_{p}\leq 1}\sum_{i\in I}|x_{i}^{\prime}(x_{i})|$ (135)
Thereby, $\Leftarrow$ and $\leq$ follow directly. To see $\Rightarrow$ and $\geq$, we choose $\tilde{x}_{i}:=x_{i}$ if $x_{i}^{\prime}(x_{i})\geq 0$ and $\tilde{x}_{i}:=-x_{i}$ otherwise. Then we find that $x_{i}^{\prime}(\tilde{x}_{i})=|x_{i}^{\prime}(x_{i})|$, and we still have $\|\tilde{x}\|_{p}=\|x\|_{p}$. Hence using $\tilde{x}$ instead of $x$ on the left hand side gives the right hand side.
In addition, note that due to
$\sum_{i\in I}x_{i}^{\prime}(x_{i})\leq\left|\sum_{i\in I}x_{i}^{\prime}(x_{i})\right|\leq\sum_{i\in I}|x_{i}^{\prime}(x_{i})|$ (136)
we might replace one of the sums by $\left|\sum_{i\in I}x_{i}^{\prime}(x_{i})\right|$ and still obtain the equivalences.
Lemma B.3.
Let $I$ be an at most countable index set and suppose that $E_{i}$ are Banach spaces for all $i\in I$. Then the mapping
${\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}\rightarrow\left({\bigoplus_{i\in I}}^{(p)}E_{i}\right)^{\prime},\quad x^{\prime}\mapsto\left(x\mapsto\sum_{i\in I}x^{\prime}_{i}(x_{i})\right)$ (137)
is an
  1. (i)
    isometric isomorphism if $p=0$ and $p^{\prime}=1$.
  2. (ii)
    isometric isomorphism if $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
  3. (iii)
    isometry if $p=\infty$ and $p^{\prime}=1$.
Proof.
A proof of (i) and (ii) can be found in [1, Theorem 16.49 and the following remark]. Let us prove (iii) here. We begin by showing that the operator is well defined in the sense that the sum converges. In fact, we even have absolute convergence, as for $x^{\prime}\in{\bigoplus_{i\in I}}^{(1)}E_{i}^{\prime}$ and $x\in{\bigoplus_{i\in I}}^{(\infty)}E_{i}$, we find that
$\sum_{i\in I}|x^{\prime}_{i}(x_{i})|\leq\sum_{i\in I}\|x_{i}^{\prime}\|_{E_{i}^{\prime}}\|x_{i}\|_{E_{i}}\leq\|x^{\prime}\|_{1}\|x\|_{\infty}<\infty\text{.}$ (138)
To show that (137) is an isometry, let $x^{\prime}\in{\bigoplus_{i\in I}}^{(1)}E_{i}^{\prime}$. We have to show that the operator norm of
${\bigoplus_{i\in I}}^{(\infty)}E_{i}\rightarrow\mathbb{R},\quad x\mapsto\sum_{i\in I}x^{\prime}_{i}(x_{i})$ (139)
agrees with $\|x^{\prime}\|_{1}$. To this end, note that
$\left\|x\mapsto\sum_{i\in I}x^{\prime}_{i}(x_{i})\right\|=\sup_{\|x\|_{\infty}\leq 1}\left|\sum_{i\in I}x^{\prime}_{i}(x_{i})\right|$ $=\sup_{\begin{subarray}{c}\|x_{i}\|_{E_{i}}\leq 1\\ \text{for all }i\in I\end{subarray}}\sum_{i\in I}x^{\prime}_{i}(x_{i})$ (140)
$=\sum_{i\in I}\sup_{\|x_{i}\|_{E_{i}}\leq 1}x^{\prime}_{i}(x_{i})$ (141)
$=\sum_{i\in I}\|x_{i}^{\prime}\|$ (142)
$=\|x^{\prime}\|_{1}\text{.}$ (143)
Concerning the absolute values, we used Remark B.2. This shows that the mapping in question is in fact an isometry. ∎
For $p$ and $p^{\prime}$ as in Lemma B.3, we introduce for $x\in{\bigoplus_{i\in I}}^{(p)}E_{i}$ and $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ the dual pairing
$\left\langle x^{\prime}\,{,}\,x\right\rangle:=\sum_{i\in I}x^{\prime}_{i}(x_{i})\text{.}$ (144)
Note that in situations (i) and (ii), Lemma B.3 allows us to describe the entire dual space. However, by part (iii) we can at least understand part of the dual space in a similar fashion as we can understand part of $(\ell^{\infty}(I))^{\prime}$ by the embedding $\iota:\ell^{1}(I)\rightarrow\ell^{1}(I)^{\prime\prime}\simeq(\ell^{\infty}(I))^{\prime}$.
Lemma B.4.
Let $I$ be an at most countable index set and suppose that $E_{i}$ are Banach spaces for all $i\in I$. Let $x^{\prime}_{i}\in E_{i}^{\prime}$ for all $i\in I$ and suppose that $1\leq p,p^{\prime}\leq\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Then the following are equivalent:
  1. (i)
    The sum
    $\sum_{i\in I}x^{\prime}_{i}(x_{i})$ (145)
    converges for all sequences $x\in\bigoplus_{i\in I}^{(p)}E_{i}$.
  2. (ii)
    The sum (145) converges absolutely for all sequences $x\in\bigoplus_{i\in I}^{(p)}E_{i}$.
  3. (iii)
    We have $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$.
Proof.
For finite $I$, all norms are finite sums, hence the assertion follows. So let us assume $I=\mathbb{N}$ without loss of generality.
“(i) $\Leftrightarrow$ (ii)” follows by Remark B.2.
“(iii) $\Rightarrow$ (i)”. Let $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$. Then an application of Lemma B.3 yields that the functional
$\gamma:{\bigoplus_{i\in I}}^{(p)}E_{i}\rightarrow\mathbb{R},\quad x\mapsto\sum_{i\in I}x^{\prime}_{i}(x_{i})$ (146)
is well-defined with $\|\gamma\|=\|x^{\prime}\|_{p^{\prime}}$. In particular, all sums in (145) converge.
“(ii) $\Rightarrow$ (iii)”. For $n\in\mathbb{N}$, let us consider the bounded operator
$S_{n}:{\bigoplus_{i\in I}}^{(p)}E_{i}\rightarrow\mathbb{R},\quad x\mapsto\sum_{i=1}^{n}x^{\prime}_{i}(x_{i})\text{.}$ (147)
Notice that for all $x\in{\bigoplus_{i\in I}}^{(p)}E_{i}$, it holds that
$\sup_{n\in\mathbb{N}}|S_{n}(x)|=\sup_{n\in\mathbb{N}}\left|\sum_{i=1}^{n}x^{\prime}_{i}(x_{i})\right|\leq\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}|x^{\prime}_{i}(x_{i})|\leq\sum_{i=1}^{\infty}|x^{\prime}_{i}(x_{i})|<\infty\text{,}$ (148)
where we used the absolute convergence of (145). Hence the uniform boundedness principle [13, Theorem 4.52] yields that $\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$.
Our next goal is to show
$\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\|_{p^{\prime}}=\|S_{n}\|\text{.}$ (149)
Case $1\leq p<\infty$. By Lemma B.3, we can understand $S_{n}\in\left(\bigoplus_{i\in I}^{(p)}E_{i}\right)^{\prime}$ as the functional associated to the sequence $(x_{i}\mathbbm{1}_{i\leq n})_{i\in I}$. Hence $\|S_{n}\|=\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\|_{p^{\prime}}$ as desired.
Case $p=\infty$. Now we have $S_{n}\in\left(\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\right)^{\prime}$. As Lemma B.3 does not directly cover this dual space, we argue differently here. To this end, note that for the operator norm we find the following:
$\|S_{n}\|=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}((x_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\leq n}^{(0)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}\text{.}$ (150)
This means that the operator norm of $S_{n}$ equals the operator norm of $S_{n}$ restricted to ${\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i}$. This restricted operator is as an element of $({\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i})^{\prime}\simeq{\bigoplus_{i\in\mathbb{N}}}^{(1)}E_{i}^{\prime}$, where we are using the isometric isomorphism from Lemma B.3. This isometric isomorphism assigns $S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}$ to the sequence $(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\in\bigoplus_{i\leq n}^{(1)}E_{i}^{\prime}$. We hence find the equality of norms
$\|S_{n}\|=\left\|S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}\right\|=\left\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\right\|_{1}\text{.}$ (151)
which finishes the second case.
Finally, we can put together the previous findings to prove that
$\|x^{\prime}\|_{p^{\prime}}=\left(\sum_{i\in\mathbb{N}}\|x^{\prime}_{i}\|^{p^{\prime}}\right)^{\frac{1}{p^{\prime}}}$ $=\sup_{n\in\mathbb{N}}\left(\sum_{i\leq n}\|x^{\prime}_{i}\|^{p^{\prime}}\right)^{\frac{1}{p^{\prime}}}$ (152)
$=\sup_{n\in\mathbb{N}}\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}\|_{p^{\prime}}$ (153)
$=\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$ (154)
as desired. ∎
Lemma B.4 does not provide concrete bounds on (145) or on $\|x^{\prime}\|_{p^{\prime}}$. However, in combination with the previous results, we can derive a quantitative version directly from Lemma B.4.
Proposition B.5.
Let $I$ be an at most countable index set and suppose that $E_{i}$ are Banach spaces for all $i\in I$. Let $x^{\prime}_{i}\in E_{i}^{\prime}$ for all $i\in I$ and suppose that $1\leq p,p^{\prime}\leq\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$. Let $C\geq 0$. Then the following are equivalent:
  1. (i)
    For all sequences $x\in{\bigoplus_{i\in I}}^{(p)}E_{i}$, the sum $\sum_{i\in I}x^{\prime}_{i}(x_{i})$ converges and it holds that
    $\left|\sum_{i\in I}x^{\prime}_{i}(x_{i})\right|\leq C\|x\|_{p}\text{.}$ (155)
  2. (ii)
    It holds that $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ with norm $\|x^{\prime}\|_{p^{\prime}}\leq C$.
Proof.
At first, let $x^{\prime}\in\bigoplus_{i\in\mathbb{N}}^{(p^{\prime})}E_{i}^{\prime}$. By the isometric embeddings from Lemma B.3, the functional
$\gamma_{x^{\prime}}:{\bigoplus_{i\in\mathbb{N}}}^{(p)}E_{i}\rightarrow\mathbb{R},\quad x\mapsto\sum_{i\in I}x_{i}^{\prime}(x_{i})$ (156)
satisfies $\|\gamma_{x^{\prime}}\|=\|x^{\prime}\|_{p^{\prime}}$.
“(i) $\Rightarrow$ (ii)”. An application of Lemma B.4 (i) $\Rightarrow$ (iii) shows that $x^{\prime}\in\bigoplus_{i\in\mathbb{N}}^{(p^{\prime})}E_{i}^{\prime}$. Hence using Remark B.2 we find that
$\|x^{\prime}\|_{p^{\prime}}=\|\gamma_{x^{\prime}}\|=\sup_{\|x\|_{p}\leq 1}\sum_{i\in I}x_{i}^{\prime}(x_{i})=\sup_{\|x\|_{p}\leq 1}\left|\sum_{i\in I}x_{i}^{\prime}(x_{i})\right|\leq C\text{.}$ (157)
“(ii) $\Rightarrow$ (i)”. The convergence of the sum follows by Lemma B.4. For $x=0$, the desired estimate is clear. For $x\neq 0$, we observe that
$\left|\frac{\sum_{i\in I}x_{i}^{\prime}(x_{i})}{\|x\|_{p}}\right|=\left|\sum_{i\in I}x_{i}^{\prime}\left(\frac{x_{i}}{\|x\|_{p}}\right)\right|\leq\sup_{\|\tilde{x}\|_{p}\leq 1}\left|\sum_{i\in I}x_{i}^{\prime}(\tilde{x}_{i})\right|=\|\gamma_{x^{\prime}}\|=\|x^{\prime}\|_{p^{\prime}}\leq C$ (158)
and note that the assertion follows by treating absolute values as explained in Remark B.2. ∎

### v2

@@
 In particular, we present a spectral decomposition of such spaces, analyse their embedding properties, and describe connections to the theory of Banach spaces of functions.
-We additionally present exemplary applications of our results to regularisation error estimation in statistical learning.
+Additionally, we present example applications of our results to regularisation error estimation in statistical learning.
+Mathematics Subject Classification (2020). Primary 46E22; Secondary 46B70, 47B34, 47G10, 47B32, 47A70, 68T05.
+Key Words. Reproducing Kernel Hilbert spaces, Interpolation spaces, Integral operators, Eigenvalues, Statistical learning theory.
 1 Introduction
-Since their introduction, reproducing kernel Hilbert spaces (RKHS) have turned out as fundamental objects in various branches of mathematical theory as well as in applications.
-As they combine the geometry of a Hilbert space with the continuous evaluation functionals of a function space, they posses many desirable properties.
+Since their introduction, reproducing kernel Hilbert spaces (RKHS) have proven to be fundamental objects in various branches of mathematical theory and in applications.
+They possess many desirable properties, as they combine the geometry of a Hilbert space with the continuous evaluation functionals of a function space.
 Apart from applications in functional analysis and numerics, statistical learning theory takes advantage of this situation by considering a large variety of kernel-based algorithms.
@@
 are important for the analysis of kernel-based least squares regression, see e.g.
+[7], for kernel interpolation, see e.g.
 [16], for estimations of the conditional mean embedding, see e.g.
-[11], and the analysis of learning curves, see e.g.
+[18], and for the analysis of learning curves, see e.g.
 [17].
-Similarly, these spaces occur naturally for certain questions related to Gaussian processes, see e.g.
-[15], [9].
+Similarly, these spaces occur naturally for certain questions related to Gaussian processes [23] [15] [28] [24], which recently have drawn interest due to their connection to neural networks [13].
+Thereby, the associated kernels are strongly related to Sobolev spaces [10], which are an important application of this paper.
 The main reason for choosing $2$ as the fine index in (1) is the spectral representation of these interpolation spaces.
-More precisely, the spaces (1) are given as the range of powers of the integral operator associated to the RKHS [18, Section 4].
-Using eigenvalues and eigenfunctions of this operator hence yields many insights into interpolation spaces and turns out as very helpful in statistical learning theory.
+More precisely, the spaces (1) are given as the range of powers of the integral operator associated to the RKHS [27, Section 4].
+Using eigenvalues and eigenfunctions of this operator hence yields many insights into interpolation spaces and is very helpful in statistical learning theory [25] [14].
 Aside from these practical advantages of fine index $2$, considering other values turned out as beneficial. In particular, the following two cases are of special interest.
-  • Fine index $1$ is of interest as it is allows for certain embedding that are not possible for other fine indexes, see e.g. [17, p. 4]
-  • Fine index $\infty$ is of interest as it allows for a precise analysis of approximation errors, see e.g. [14, Theorem 3.1].
+  • In certain boundary cases, interpolation spaces of fine index $1$ enable embeddings into $\mathrm{L}^{\infty}(\nu)$ that are not possible for other fine indexes.
+For examples and applications of such embeddings, see e.g.
+[26, p.
+4].
+  • Interpolation spaces of fine index $\infty$ allow for a precise analysis of approximation errors, see e.g. [22, Theorem 3.1].
 For these reasons, we consider real interpolation spaces of the general form
@@
 Our aim is to present methods and results that treat such interpolation spaces for all fine indexes $r$ in a joint fashion.
-The results, we will derive address the following main questions:
+The results we will derive address the following main questions:
   • Do the spaces (2) and their norms still possess a spectral representation?
-  • The interpolation spaces consist of equivalence classes of functions at first. When can we identify the spaces (2) with Banach spaces consisting of functions?
-  • When are the spaces (2) continuously embedded into $\mathrm{L}^{\infty}$?
-The first question is clearly motivated by the practicality of spectral representations, as indicated by the papers mentioned earlier.
+  • The interpolation spaces consist of equivalence classes of functions. When can we identify the spaces (2) with Banach spaces consisting of functions?
+  • When are the spaces (2) continuously embedded into $\mathrm{L}^{\infty}(\nu)$?
+The first question is clearly motivated by the practicality of spectral representations, as indicated by the papers mentioned above.
 Positive answers to the second question have applications in all situations where a point evaluation of functions is necessary.
 A main example is given by the theory of Gaussian processes in the context of machine learning, as paths cannot be equivalence classes but need to be defined everywhere, see e.g.
-[12, Assumption 1].
+[20, Assumption 1], or by the analysis of function spaces generated by neural networks, see e.g.
+[2].
 The third question will turn out to be tightly connected to the second one, but applications go beyond that.
 In particular, the statistical analysis of kernel-based learning algorithms relies on such embeddings, see e.g.
-[16, Theorem 7.23] [5, Section 3] [17, p.
-4].
+[25, Theorem 7.23] [7, Section 3] [26, p.
+4] [19, Section 4].
+Our results on the second and third question improve earlier findings such as [23, Section 5] [27, Section 5] even in the Hilbert space case $r=2$.
 This work is organised as follows: we start by introducing the necessary notation and fundamental definitions in Section 2.
 Then the main results are provided in Section 3.
-We present an exemplary application to error estimation in statistical learning theory in Section 4.
+In Section 4, we present an example application to error estimation in statistical learning theory.
 All proofs and the ideas behind the results will be given in Section 5.
-Finally, some auxiliary results can be found in two appendices.
+Finally, some auxiliary results can be found in the three appendices.
 2 Preliminaries
-Given normed $\mathbb{R}$-vector spaces $E_{1}$ and $E_{2}$, we use $E_{1}\simeq E_{2}$ to denote isometric isomorphy, $E_{1}\triangleq E_{2}$ for norm equivalent spaces and $E_{1}\hookrightarrow E_{2}$ if $E_{1}\subset E_{2}$ and the embedding is continuous.
+Given normed $\mathbb{R}$-vector spaces $E_{1}$ and $E_{2}$, we use $E_{1}\simeq E_{2}$ to denote isometric isomorphism, $E_{1}\triangleq E_{2}$ for norm equivalent spaces and $E_{1}\hookrightarrow E_{2}$ if $E_{1}\subset E_{2}$ and the embedding is continuous.
 For a measurable space $(X,\mathcal{A})$, we denote the set of measurable functions $X\rightarrow\mathbb{R}$ by $\mathcal{L}^{0}(X)$.
-The set of bounded functions in $\mathcal{L}^{0}(X)$ is denoted by $\mathcal{L}^{\infty}(X)$ and this space is equipped with the supremum norm.
+The set of bounded functions in $\mathcal{L}^{0}(X)$ is denoted by $\mathcal{L}^{\infty}(X)$, and this space is equipped with the supremum norm.
 Let us assume that additionally, a measure $\nu$ is given on $X$.
@@
 For more details, see [3, Chapter 5] or [4, Chapter 3].
-In the following, we only consider interpolation spaces of Banach spaces $E_{1}$ and $E_{2}$ where $E_{1}\hookrightarrow E_{2}$.
-To define the real interpolation spaces, we use the so-called $K$-method. To this end, for all $x\in E_{2}$ and $t>0$, we define the $K$-functional as
-$K(x,t,E_{2},E_{1}):=\inf_{y\in E_{1}}\left(\|x-y\|_{E_{2}}+t\|y\|_{E_{1}}\right)\text{.}$ (3)
-For $x\in E_{2}$ and the parameters $0<\theta<1$ and $1\leq r<\infty$, this enables us to set
-$\|x\|_{[E_{2},E_{1}]_{\theta,r}}:=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,E_{2},E_{1})\mathrm{d}t\right)^{\frac{1}{r}}$ (4)
+In the following, we only consider interpolation spaces of Banach spaces $E_{1}$ and $E_{2}$ where $E_{2}\hookrightarrow E_{1}$.
+To define the real interpolation spaces, we use the so-called $K$-method. To this end, for all $x\in E_{1}$ and $t>0$, we define the $K$-functional as
+$K(x,t,E_{1},E_{2}):=\inf_{y\in E_{2}}\left(\|x-y\|_{E_{1}}+t\|y\|_{E_{2}}\right)\text{.}$ (3)
+For $x\in E_{1}$ and the parameters $0<\theta<1$ and $1\leq r<\infty$, this enables us to set
+$\|x\|_{[E_{1},E_{2}]_{\theta,r}}:=\left(\int_{0}^{\infty}t^{-\theta r-1}K^{r}(x,t,E_{1},E_{2})\mathrm{d}t\right)^{\frac{1}{r}}$ (4)
 and similarly, for $r=\infty$, we let
-$\|x\|_{[E_{2},E_{1}]_{\theta,\infty}}:=\sup_{t>0}\left(t^{-\theta}K(x,t,E_{2},E_{1})\right)\text{.}$ (5)
+$\|x\|_{[E_{1},E_{2}]_{\theta,\infty}}:=\sup_{t>0}\left(t^{-\theta}K(x,t,E_{1},E_{2})\right)\text{.}$ (5)
 Now the interpolation space of $E_{1}$ and $E_{2}$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined as
-$[E_{2},E_{1}]_{\theta,r}:=\{x\in E_{2}\,:\,\|x\|_{[E_{2},E_{1}]_{\theta,r}}<\infty\}$ (6)
-and this space is equipped with the $\|\,\cdot\,\|_{[E_{2},E_{1}]_{\theta,r}}$-norm.
+$[E_{1},E_{2}]_{\theta,r}:=\{x\in E_{1}\,:\,\|x\|_{[E_{1},E_{2}]_{\theta,r}}<\infty\}$ (6)
+and this space is equipped with the $\|\,\cdot\,\|_{[E_{1},E_{2}]_{\theta,r}}$-norm.
 These spaces are again Banach spaces, see [3, Proposition 1.8 in Chapter 5].
-In fact, such interpolation spaces perform a kind of interpolation of Banach spaces, as they are “in between” $E_{1}$ and $E_{2}$ in the sense that $E_{1}\hookrightarrow[E_{2},E_{1}]_{\theta,r}\hookrightarrow E_{1}$, see Lemma A.2.
+In fact, such interpolation spaces perform a kind of interpolation of Banach spaces, as they are “in between” $E_{1}$ and $E_{2}$ in the sense that $E_{2}\hookrightarrow[E_{1},E_{2}]_{\theta,r}\hookrightarrow E_{1}$, see Lemma A.2.
 Note that the definition of real interpolation spaces presented here is just one of several alternatives.
@@
 are bounded functionals on $E$, then $E$ is called a Banach space of functions (BSF), see [29] for more details.
-These spaces recently attracted interest, as they yield insights into neural networks [2].
 If additionally, the norm of such a Banach space of functions $H$ is induced by an inner product, for each $x\in X$, we can find a function $k(\cdot,x)\in H$ that is associated to $\delta_{x}$ by the Riesz representation theorem.
 Then $H$ is a reproducing kernel Hilbert space (RKHS) on $X$ with kernel $k$.
-We refer to [16, Chapter 4] for fundamental results on RKHS.
+We refer to [25, Chapter 4] for fundamental results on RKHS.
 Now let $(X,\mathcal{A},\mu)$ be a measure space.
@@
 see e.g.
-[18, Lemma 2.2].
+[27, Lemma 2.2].
 Let us denote the non-zero eigenvalues of $T_{k}$ in non-increasing order by $\mu=(\mu_{i})_{i\in I}$, where $I=\{1,...,n\}$ or $I=\mathbb{N}$.
 Due to the positivity and compactness of $T_{k}$, we know that $\mu_{i}>0$ and, in case that $I$ is infinite, $\lim_{i\rightarrow\infty}\mu_{i}=0$.
-As proven in [18, Lemma 2.12], in such a situation, there are functions $e_{i}\in H$ such that $([e_{i}]_{\sim})_{i\in I}\subset\mathrm{L}^{2}(\nu)$ forms an ONS in $\mathrm{L}^{2}(\nu)$ consisting of eigenvectors of the integral operator $T_{k}$, such that additionally $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ forms an ONS inside $H$.
+As proven in [27, Lemma 2.12], in such a situation, there are functions $e_{i}\in H$ such that $([e_{i}]_{\sim})_{i\in I}\subset\mathrm{L}^{2}(\nu)$ forms an ONS in $\mathrm{L}^{2}(\nu)$ consisting of eigenvectors of the integral operator $T_{k}$, such that additionally, $(\sqrt{\mu_{i}}e_{i})_{i\in I}$ forms an ONS inside $H$.
 The range of $I_{k}$, that is $H]_{\sim}$, becomes a Hilbert space by using the quotient norm $\|[f]_{\sim}\|:=\inf\{\|g\|_{H}\,|\,I_{k}g=[f]_{\sim}\}$.
@@
 Assumption H.
-Suppose that $(X,\mathcal{A},\mu)$ is a measure space and that $H$ is an RKHS of a measurable kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$.
+Suppose that $(X,\mathcal{A},\nu)$ is a measure space and that $H$ is an RKHS of a measurable kernel $k$ on $X$ such that $H$ is compactly embedded into $\mathrm{L}^{2}(\nu)$.
 Then, for a suitable index set $I=\{1,...,n\}$ or $I=\mathbb{N}$, we denote the eigenvalues of $T_{k}:\mathrm{L}^{2}(\nu)\rightarrow\mathrm{L}^{2}(\nu)$ by $\mu_{i}$ and the associated eigenfunctions by $e_{i}$ for $i\in I$.
-Assumption H+. Let Assumption H be satisfied. In addition, let the measure space $(X,\mathcal{A},\mu)$ be complete and $\mu$ be $\sigma$-finite. Furthermore, assume that $k$ is a bounded kernel.
+Assumption H+. Let Assumption H be satisfied. In addition, let $\nu$ be $\sigma$-finite. Furthermore, assume that $k$ is a bounded kernel.
+Let us recall two concepts from topology and measure theory.
+Firstly, let $X$ be a set and suppose that $M$ is a set of functions $X\rightarrow\mathbb{R}$.
+Then the coarsest topology $\tau(M)$ on $X$ such that all functions in $M$ are continuous is called the initial topology on $X$ with respect to $M$.
+Secondly, let $(X,\mathcal{A})$ be a measurable space and $(X,\mathcal{T})$ a topological space.
+Then a measure $\nu$ on $(X,\mathcal{A})$ is called $\mathcal{T}$-positive, if $\mathcal{T}\subset\mathcal{A}$, i.e.
+each open set is measurable, and if for all non-empty open sets $A$, it holds that $\nu(A)>0$.
+Assumption T.
+Let $(X,\mathcal{A},\nu)$ be a measure space such that $\mathrm{L}^{2}(\nu)$ is separable.
+Moreover, let $H$ be an RKHS over $X$ of a measurable kernel $k$ such that Assumption H is satisfied and $\nu$ is $\tau(H)$-positive.
+We use $I$, $e_{i}$ and $\mu_{i}$ as in Assumption H.
+Remark 2.1.
+We give a sufficient condition that allows to verify Assumption T in the most common situations. To this end, suppose a topology $\mathcal{T}$ on $X$ is given such that
+  1. (i)
+    $\mathcal{T}\subset\mathcal{A}$, i.e. $\mathcal{A}$ contains the Borel-$\sigma$-algebra of $\mathcal{T}$,
+  2. (ii)
+    $\nu$ is $\mathcal{T}$-positive, and
+  3. (iii)
+    $k$ is $\mathcal{T}\otimes\mathcal{T}$-continuous.
+Then, due to (iii), all functions in $H$ are $\mathcal{T}$-continuous and hence $\tau(H)\subset\mathcal{T}$.
+By (ii), it follows that all non-empty open sets in $\tau(H)$ have positive measure, so $\nu$ is $\tau(H)$-positive as well.
+Hence $\tau(H)\subset\mathcal{A}$ is $\nu$-positive.
+Note however that the separability of $\mathrm{L}^{2}(\nu)$ has to be checked separately.
+In particular, assume that $X$ is a subset of $\mathbb{R}^{n}$ that satisfies $\mathrm{clos}(\mathrm{int}(X))=\mathrm{clos}(X)$, where $\mathrm{clos}$ and $\mathrm{int}$ denote the closure and the interior, respectively.
+Then Assumption T is satisfied if we use the Borel-$\sigma$-algebra and the Lebesgue measure with a continuous kernel on $X$.
+More generally, we may replace $\mathbb{R}^{n}$ by a second-countable locally compact group equipped with the Borel-$\sigma$-algebra and the Haar measure.
 Power spaces
@@
 in terms of eigenvalues and eigenfunctions of the operator $T_{k}$.
-To this end, for a “weight function” $w:I\rightarrow(0,\infty)$, we consider the following partition of $I$:
-$M_{j}:=\left\{i\in I\,\middle|\,2^{j}<w(i)\leq 2^{j+1}\right\},\quad j\in\mathbb{Z}\text{.}$ (9)
+To this end, for a “weight sequence” $\mu:I\rightarrow(0,\infty)$, we consider the following partition of $I$:
+$M_{j}:=\left\{i\in I\,\middle|\,2^{j}<\mu_{i}\leq 2^{j+1}\right\},\quad j\in\mathbb{Z}\text{.}$ (9)
 Definition 2.2.
@@
 Note that in the case $r=2$, one finds that $\ell^{(\mu,\theta,r)}=\ell^{2}(\mu^{-\theta})$ with equal norms.
-In consequence, the following definition generalises the notion of a power of an RKHS from [18, Section 4].
+In consequence, the following definition generalises the notion of a power of an RKHS from [27, Section 4].
 Definition 2.3.
-Let $H$, $e_{i}$ and $\mu_{i}$ be as introduced in Assumption H. Then the power space of $H$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined by
-$H_{\nu}^{(\theta,r)}:=\left\{f\in\mathrm{L}^{2}(\nu)\,\middle|\,\exists b\in\ell^{(\mu,\theta,r)}\text{ with }f=\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\}$ (12)
+Let $H$, $e_{i}$ and $\mu_{i}$ be as introduced in Assumption H. Then the power space of $H$ to the parameters $0<\theta<1$ and $1\leq r\leq\infty$ is defined as
+$[H]_{\sim}^{(\theta,r)}:=\left\{f\in\mathrm{L}^{2}(\nu)\,\middle|\,\exists b\in\ell^{(\mu,\theta,r)}\text{ with }f=\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\}$ (12)
 with norm given by
-$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}:=\|b\|_{(\mu,\theta,r)}\text{.}$ (13)
-Note that this norm is well defined due to the uniqueness of the Fourier coefficients $b_{i}$.
-Recall that in [18, Section 4] only the spaces $H_{\nu}^{\theta,2}$ were introduced and investigated.
-Moreover, for $r\neq 2$, the spaces $H_{\nu}^{(\theta,r)}$ are no longer Hilbert spaces.
+$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}:=\|b\|_{(\mu,\theta,r)}\text{.}$ (13)
+Note that this norm is well-defined due to the uniqueness of the Fourier coefficients $b_{i}$.
+Recall that in [27, Section 4], only the spaces $H]_{\sim}^{(\theta,2)}$ were introduced and investigated.
+Moreover, for $r\neq 2$, the spaces $H]_{\sim}^{(\theta,r)}$ are no longer Hilbert spaces.
 3 Results
-Our three main results are provided in the following.
-The first result states the norm equivalence of the power spaces $H_{\nu}^{\theta,r}$ to the interpolation spaces of the form (2), as previously mentioned.
-This is the foundation for the subsequent result, which shows that we can view these interpolation spaces as a BSF, in case they are embedded into $\mathrm{L}^{\infty}(\nu)$.
-The third result then analyses in which situations such an embedding exists.
+Our main results are provided in the following.
+The first result states the norm equivalence of the power spaces $H]_{\sim}^{\theta,r}$ to the interpolation spaces of the form (2), as previously mentioned.
+This is the foundation for the subsequent two results, which show that we can view these interpolation spaces as BSFs, in case they are embedded into $\mathrm{L}^{\infty}(\nu)$.
+The last result then analyses in which situations such an embedding exists.
 Theorem 3.1.
-Let $0<\theta<1$ and $1\leq r\leq\infty$ and suppose that $H$ is as in Assumption H. Then
-$H_{\nu}^{(\theta,r)}=[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ (14)
+Let $0<\theta<1$, $1\leq r\leq\infty$, and suppose that $H$ is as in Assumption H. Then
+$[H]_{\sim}^{(\theta,r)}=[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}$ (14)
 and the norms on these spaces are equivalent.
@@
 Let $(X,\mathcal{A},\nu)$ and $H$ be as in Assumption H+.
-Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous.
+Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous.
 Then there is a BSF $\overline{H}_{\nu}^{(\theta,r)}$ on $X$ such that the operator
-$\overline{H}_{\nu}^{(\theta,r)}\rightarrow H_{\nu}^{(\theta,r)},\quad f\mapsto[f]_{\sim}$ (15)
+$\overline{H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)},\quad f\mapsto[f]_{\sim}$ (15)
 is a well-defined isometric isomorphism with
-$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (16)
-The existence of an embedding of the power space $H_{\nu}^{(\theta,r)}$ into $\mathrm{L}^{\infty}(\nu)$ seems to be a crucial requirement of Theorem 3.2.
+$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (16)
+Additionally, there is a $\nu$-null set $N\subset X$ such that for all $f\in\overline{H}_{\nu}^{(\theta,r)}$, there is a $(b_{i})_{i\in I}\in\ell^{(\mu,\theta,r)}$ such that $f(x)=\sum_{i\in I}b_{i}e_{i}(x)$ for all $x\in X\setminus N$.
+Theorem 3.2 ensures the existence of a BSF, whose functions may differ from the actual functions in $H$ on sets of measure zero.
+Under Assumption T, we can make the BSF more explicit.
+Then we can take orthogonal series of eigenfunctions as elements of the BSF, which finally enables to embed $H$ into this BSF, as the following Theorem states.
+Theorem 3.3.
+Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption T.
+Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous.
+Then the following statements are true:
+  1. (i)
+    The series $\sum_{i\in I}b_{i}e_{i}$ converges uniformly for all $b\in\ell^{(\mu,\theta,r)}$.
+  2. (ii)
+    The space
+    ${H}_{\nu}^{(\theta,r)}:=\left\{\sum_{i\in I}b_{i}e_{i}\,\middle|\,b\in\ell^{(\mu,\theta,r)}\right\}$ (17)
+    is a BSF, where the BSF-norm is inherited from the $\ell^{(\mu,\theta,r)}$-norm of the Fourier coefficients $b_{i}$.
+  3. (iii)
+    The mapping
+    ${H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)},\quad\sum_{i\in I}b_{i}e_{i}\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ (18)
+    is an isometric isomorphism.
+  4. (iv)
+    It holds that
+    $\|{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{C}_{b}(X,\tau(H))\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (19)
+  5. (v)
+    We have $H\hookrightarrow{H}_{\nu}^{(\theta,r)}$.
+The existence of an embedding of the power space $H]_{\sim}^{(\theta,r)}$ into $\mathrm{L}^{\infty}(\nu)$ seems to be a crucial requirement of Theorems 3.2 and 3.3.
 Let us elaborate in which situations we can expect such an embedding.
-For fine-index $r=2$, this question has been approached in [18, Section 5].
+For fine index $r=2$, this question has been approached in [27, Section 5].
 We extend this approach to $1\leq r\leq\infty$.
 Theorem 3.4.
-Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+.
-Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$, where $\frac{1}{r}+\frac{1}{r^{\prime}}=1$.
-Then the following statements are equivalent:
+Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H.
+Moreover, let $\kappa>0$, $0<\theta<1$, and $1\leq r\leq\infty$, where $\frac{1}{r}+\frac{1}{r^{\prime}}=1$.
+Consider the following statements:
   1. (i)
     It holds that
-    $\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa$ (17)
+    $\sup_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa\text{.}$ (20)
   2. (ii)
-    The embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous with operator norm at most $\kappa$.
+    It holds that
+    $\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa\text{.}$ (21)
+  3. (iii)
+    The embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous with operator norm at most $\kappa$.
+Then we always have (i) $\Rightarrow$ (ii) $\Rightarrow$ (iii).
+Moreover, under Assumption H+, we have (i) $\Rightarrow$ (ii) $\Leftrightarrow$ (iii), and under Assumption T, we even have (i) $\Leftrightarrow$ (ii) $\Leftrightarrow$ (iii).
+Remark 3.5.
+Suppose that Assumption H is satisfied. Recall that the power kernel associated to $k$ is defined by its Mercer representation as
+$k_{\nu}^{\theta}(x,x^{\prime})=\sum_{i\in I}\mu_{i}^{\theta}e_{i}(x)e_{i}(x^{\prime})$ (22)
+for $x,x^{\prime}\in X$, see [27, Proposition 4.2]. Using this notion, we can reformulate the terms in (20) and (21) as follows.
+$\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}=\left\|\sum_{i\in I}e_{i}(x)\mu_{i}^{\theta}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}$ $=\left\|\left[\sum_{i\in I}e_{i}(x)\mu_{i}^{\theta}e_{i}\right]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}$ (23)
+$=\left\|\left[k_{\nu}^{\theta}(x,\cdot)\right]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r^{\prime})}}\text{.}$ (24)
+Moreover, if $\sum_{i\in I}b_{i}e_{i}$ converges pointwise everywhere for all $b\in\ell^{(\mu,\theta,r^{\prime})}$, we may define $H_{\nu}^{(\theta,r^{\prime})}$ as in (17) and remove all $\cdot\,$$_{\sim}$ to consider functions instead of equivalence classes.
+Under Assumption T, this is possible due to Theorem 3.3.
+Remark 3.6.
+Suppose there is a $C>0$ such that the eigenfunctions $e_{i}$ satisfy $\|e_{i}\|_{\mathrm{L}^{\infty}(\nu)}<C$ for all $i\in I$.
+Then, by a countable union of null sets, we find that $|e_{i}(x)|<C$ for all $i\in I$ and for $\nu$-almost all $x\in X$.
+Then
+$\operatorname*{\mathrm{ess-sup}}_{x\in X}\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq C\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}\text{,}$ (25)
+and hence condition (21) is satisfied for some $\kappa>0$ whenever $\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}<\infty$.
+More explicitly, the term of interest $\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}$ is given by
+$\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{r^{\prime}}{2}}\right)^{\frac{1}{r^{\prime}}}\quad\text{if }1\leq r^{\prime}<\infty\quad\text{ and by }\quad\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{1}{2}}\quad\text{if }r^{\prime}=\infty\text{.}$ (26)
+Remark 3.7.
+We illustrate the statements in Theorem 3.4 in a concrete situation.
+To this end, suppose that $\mu_{i}=i^{-\alpha}$ for $\alpha>1$, which is linked to Sobolev spaces [11].
+Then we find for $j\in\mathbb{Z}$ that
+$M_{j}=\{i\in\mathbb{N}\,|\,2^{j}<i^{-\alpha}\leq 2^{j+1}\}=\{i\in\mathbb{N}\,|\,2^{-\frac{j+1}{\alpha}}\leq i<2^{-\frac{j}{\alpha}}\}=[2^{-\frac{j+1}{\alpha}},2^{-\frac{j}{\alpha}})\cap\mathbb{N}\text{,}$ (27)
+so $M_{j}=\{\}$ for $j\geq 0$. We hence note that $\min M_{j}\geq 2^{-\frac{j+1}{\alpha}}$ and $\max M_{j}\leq 2^{-\frac{j}{\alpha}}$. Furthermore, we have
+$2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}-1\leq|M_{j}|\leq 2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}+1$ (28)
+By that, we estimate for $j<0$
+$\sum_{i\in M_{j}}i^{-1}\leq|M_{j}|2^{\frac{j+1}{\alpha}}\leq(2^{-\frac{j}{\alpha}}-2^{-\frac{j+1}{\alpha}}+1)2^{\frac{j+1}{\alpha}}\leq 2^{\frac{1}{\alpha}}-1+2^{\frac{j+1}{\alpha}}\leq 1+2^{\frac{j+1}{\alpha}}\leq 1+2^{\frac{1}{\alpha}}\text{.}$ (29)
+Hence for $\theta=\frac{1}{\alpha}$ with $r=1$ and $r^{\prime}=\infty$, we get
+$\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}=\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{1}{2}}=\sup_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}i^{-1}\right)^{\frac{1}{2}}\leq\left(1+2^{\frac{1}{\alpha}}\right)^{\frac{1}{2}}<\infty\text{.}$ (30)
+However, for $r=2$ with $r^{\prime}=2$, we get
+$\|(\mu_{i}^{\theta})_{i\in I}\|_{(\mu,\theta,r^{\prime})}=\left(\sum_{j\in\mathbb{Z}}\left(\sum_{i\in M_{j}}\mu_{i}^{\theta}\right)^{\frac{r^{\prime}}{2}}\right)^{\frac{1}{r^{\prime}}}=\left(\sum_{j\in\mathbb{Z}}\sum_{i\in M_{j}}i^{-1}\right)^{\frac{1}{2}}=\left(\sum_{i\in\mathbb{N}}i^{-1}\right)^{\frac{1}{2}}=\infty\text{.}$ (31)
+We see that in this situation, the fine index $r$ decides whether the conditions (20) and (21) are met.
 4 Application to statistical learning
@@
 Suppose there is a probability measure $P$ on $X\times Y$, wherefrom only some samples $(x_{1},y_{1}),...,(x_{n},y_{n})$ are known.
-Typically, one searches for a prediction function $f:X\rightarrow\mathbb{R}$ that predicts an $Y$-value associated to a given $X$-value with low risk:
+Typically, one searches for a prediction function $f:X\rightarrow\mathbb{R}$ that predicts an $Y$-value associated to a given $X$-value with low risk
 $\mathcal{R}_{L,P}(f):=\int_{X\times Y}L(y,f(x))dP(x,y)\text{.}$ (32)
@@
 $f_{P,\lambda}\in\mathrm{argmin}_{f\in F}\;\lambda\|f\|_{H}^{2}+\mathcal{R}_{L,P}(f)$ (33)
-where $P$ is in practice replaced by the empirical probability measure assigned to the given data. As done in [6], we assume in this section that the kernel of $H$ is bound by $1$.
-Assuming that $L$ is the least squares loss, $f_{P,\lambda}$ can alternatively be found by the eigen-decomposition of $H$, see [18, Theorem 5.4]:
+where $P$ is in practice replaced by the empirical probability measure assigned to the given data. As done in [8], we assume in this section that the kernel of $H$ is bounded by $1$.
+Assuming that $L$ is the least squares loss, $f_{P,\lambda}$ can alternatively be found by the eigen-decomposition of $H$, see [27, Theorem 5.4]:
 $f_{P,\lambda}=\sum_{i\in I}\frac{\mu_{i}}{\lambda+\mu_{i}}\left\langle f_{L,P}^{*}\,{,}\,[e_{i}]_{\sim}\right\rangle_{\mathrm{L}^{2}(P_{X})}e_{i}$ (34)
 where we assume that $P_{X}$ is the marginal distribution, that Y=[1,1]Y=[-1,1] and that $H$, the eigenvalues $\mu_{i}$ and the basis $e_{i}$ are as in Assumption H.
-By this reformulation, we can to describe various other methods of regularised minimisation than (19) by replacing the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ by a more general function $\mu_{i}g_{\lambda}(\mu_{i})$.
+By this reformulation, we can describe various other methods of regularised minimisation than (33) by replacing the term $\frac{\mu_{i}}{\lambda+\mu_{i}}$ with a more general function $\mu_{i}g_{\lambda}(\mu_{i})$.
 The following conditions taken from [8] ensure that a reasonable algorithm emerges.
@@
 Let $0<\beta<1$, $0<\theta<1$ and $1\leq r\leq\infty$.
-If $f_{L,P}^{*}\in[\mathrm{L}^{2}(P_{X}),[H]_{\sim}]_{\beta,r}$, the following holds.
+If $f_{L,P}^{*}\in[\mathrm{L}^{2}(P_{X}),[H]_{\sim}]_{\beta,r}$, the following holds:
   1. (i)
@@
 Such estimates are play an important role in the statistical analysis of kernel-based learning algorithms, see e.g.
-[16, Theorem 7.23].
+[25, Theorem 7.23].
 Note that the required existence of the embedding to $\mathrm{L}^{\infty}(\nu)$ may be checked via Theorem 3.4.
@@
     $\leq\lambda^{\gamma}\frac{\left(\frac{\gamma}{1-\gamma}\right)^{\gamma}}{1+\frac{\gamma}{1-\gamma}}$ (46)
-    $=\lambda^{\gamma}\gamma^{\gamma}(1-\gamma)^{1-\gamma}$ (33)
+    $=\lambda^{\gamma}\gamma^{\gamma}(1-\gamma)^{1-\gamma}\text{,}$ (47)
     which can be seen by finding the zero of the derivative of the expression in question. Hence for the estimate in Proposition 4.2, we get
-    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\lambda^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}$ (34)
+    $\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}\leq C_{3}\lambda^{\gamma}\|f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\beta,r}}\text{,}$ (48)
     where we additionally use that $\gamma^{\gamma}(1-\gamma)^{1-\gamma}\leq 1$.
@@
     Landweber iteration. We have $g_{t}(\mu)=\sum_{i=0}^{t-1}(1-\mu)^{i}$ with $\lambda=t^{-1}$. For $\mu<1$ we reformulate this by the geometric sum as $g_{t}=\frac{1-(1-\mu)^{t}}{\mu}$ to find that
-    $\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\gamma}\right)=\sup_{i\in I}(1-\mu_{i})^{t}\mu_{i}^{\gamma}\leq\sup_{0<\mu<1}(1-\mu)^{t}\mu^{\gamma}=\left(\frac{t}{\gamma+t}\right)^{t}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}$ (35)
+    $\sup_{i\in I}\left(\big|1-\mu_{i}g_{\lambda}(\mu_{i})\big|\mu_{i}^{\gamma}\right)=\sup_{i\in I}(1-\mu_{i})^{t}\mu_{i}^{\gamma}\leq\sup_{0<\mu<1}(1-\mu)^{t}\mu^{\gamma}=\left(\frac{t}{\gamma+t}\right)^{t}\left(\frac{\gamma}{\gamma+t}\right)^{\gamma}\text{,}$ (49)
     which can be seen by finding the zero of the derivative of the expression in question. Hence for the estimate in Proposition 4.2, we get
@@
 Theorem 5.1.
-Let $(X,\mathcal{A},\nu)$ be a measure space, and $w:X\rightarrow[\epsilon,\infty)$ be measurable, where $\epsilon>0$ is some constant.
-Furthermore, let $s>1$, $\theta\in(0,1)$, and $r\in[1,\infty)$.
+Let $(X,\mathcal{A},\nu)$ be a measure space, and $w:X\rightarrow[\varepsilon,\infty)$ be measurable, where $\varepsilon>0$ is some constant.
+Furthermore, let $s>1$, $0<\theta<1$, and $1\leq r<\infty$.
 Then
@@
 Proof.
-The assertion directly follows from [7, Theorem 3.3], where we set $\mathscr{X}$ as $\mathrm{L}^{p}(\nu)$ and used the weight $\omega=w^{\frac{1}{p}}$, as the author applies the weight to the functions instead of the measures as we do here.
+The assertion directly follows from [9, Theorem 3.3], where we set $\mathscr{X}$ as $\mathrm{L}^{p}(\nu)$ and use the weight $\omega=w^{\frac{1}{p}}$, as the author applies the weight to the functions instead of the measures as we do here.
 Finally, we need to recall that the $K$-method and $J$-method used in [9] yield equivalent norms on the interpolation space.
@@
 To this end, we generalise the definition of $M_{j}$ from (9) to a general measure space $(X,\mathcal{A},\nu)$ and a measurable “weight” function $w:X\rightarrow(0,\infty)$:
-$M_{j}:=\left\{x\in X\,\middle|\,\frac{1}{2^{j+1}}\leq\frac{1}{w(x)}<\frac{1}{2^{j}}\right\}\text{.}$ (38)
+$M_{j}:=\left\{x\in X\,\middle|\,\frac{1}{2^{j+1}}\leq\frac{1}{w(x)}<\frac{1}{2^{j}}\right\}$ (52)
 Lemma 5.2.
-Let $(X,\mathcal{A},\nu)$ be a measure space and let $\varepsilon>0$, and $w:X\rightarrow[\epsilon,\infty)$ be measurable. Furthermore, let $\theta\in(0,1)$, $r\in[1,\infty)$, and let
+Let $(X,\mathcal{A},\nu)$ be a measure space, let $\varepsilon>0$, and $w:X\rightarrow[\varepsilon,\infty)$ be measurable. Furthermore, let $0<\theta<1$, $1\leq r<\infty$, and let
 $c_{j}(f):=\left(\int_{M_{j}}|f|^{p}w^{\theta}\mathrm{d}\nu\right)^{\frac{1}{p}}$ (53)
@@
 We write $H_{I}:=\overline{\operatorname{\mathrm{span}}\{\sqrt{\mu_{i}}e_{i}|i\in I\}}$ for the subspace spanned by this ONS.
-By [18, Formula (17)], we know that $H_{I}=\overline{\text{Im}I_{k}^{*}}$, and hence [18, Formula (19)] shows $H_{I}^{\perp}=\text{Ker}I_{k}$.
+By [27, Formula (17)], we know that $H_{I}=\overline{\text{Im}I_{k}^{*}}$, and hence [27, Formula (19)] shows $H_{I}^{\perp}=\text{Ker}I_{k}$.
 If we choose an ONB $(\tilde{e}_{j})_{j}$ of $H_{I}^{\perp}$ on a new index set $J$, we thus find
-$[H]_{\sim}$ $=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}+\sum_{j\in J}b_{i}\tilde{e}_{j}\right]_{\sim}\,\middle|\,(a_{i})\in\ell^{2}(I),(b_{j})\in\ell^{2}(J)\right\}$ (44)
-$=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]\,\middle|\,(a_{i})\in\ell^{2}(I)\right\}$ (45)
+$[H]_{\sim}$ $=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}+\sum_{j\in J}b_{j}\tilde{e}_{j}\right]_{\sim}\,\middle|\,a\in\ell^{2}(I),b\in\ell^{2}(J)\right\}$ (58)
+$=\left\{\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]\,\middle|\,a\in\ell^{2}(I)\right\}$ (59)
 $=[H_{I}]_{\sim}\text{.}$ (60)
@@
 $\sum_{i\in I}b_{i}[e_{i}]_{\sim}=[h]_{\sim}=\left[\sum_{i\in I}a_{i}\sqrt{\mu_{i}}e_{i}\right]_{\sim}=\sum_{i\in I}a_{i}\sqrt{\mu_{i}}[e_{i}]_{\sim}\text{,}$ (67)
-and thus $a_{i}\sqrt{\mu_{i}}=b_{i}$. Since $(a_{i})\in\ell^{2}(I)$, we get
+and thus $a_{i}\sqrt{\mu_{i}}=b_{i}$. Since $a\in\ell^{2}(I)$, we get
 $\|b\|_{\ell^{2}(\mu^{-1})}^{2}=\sum_{i\in I}b_{i}^{2}\mu_{i}^{-1}=\sum_{i\in I}\left(\frac{b_{i}}{\sqrt{\mu_{i}}}\right)^{2}=\sum_{i\in I}a_{i}^{2}=\|h\|_{H_{I}}^{2}\text{.}$ (68)
@@
 It maps functions to their $\mathrm{L}^{2}(\nu)$-Fourier coefficients with respect to the ONS $e_{i}$.
-The map $\beta:\ell^{(\mu,\theta,r)}\rightarrow H_{\nu}^{(\theta,r)}$ given by $b\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ is an isometric isomorphism by Definition 2.2.
+The map $\beta:\ell^{(\mu,\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)}$ given by $b\mapsto\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ is an isometric isomorphism by Definition 2.3.
 We note that for $f\in[\mathrm{L}^{2}(\nu),[H]_{\sim}]_{\theta,r}$, we have
@@
 By Lemma A.2, we know that $\mathrm{L}^{2}(\nu),[H]_{\sim}]_{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$.
-Hence we can apply Theorem 3.1 to see that for the power spaces, we have $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$ as well.
+Hence we can apply Theorem 3.1 to see that for the power spaces, we have $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{2}(\nu)$ as well.
 However, the operator norms of these two embedding operators might differ by a constant, as the norms are equivalent, but not necessarily equal.
 Lemma 5.6.
-Let $0<\theta<1$ and $1\leq r\leq\infty$. If $b\in\ell^{(\mu,\theta,r)}$, then $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in the power space $H_{\nu}^{(\theta,r)}$.
+Let $0<\theta<1$ and $1\leq r\leq\infty$. If $b\in\ell^{(\mu,\theta,r)}$, then $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in the power space $H]_{\sim}^{(\theta,r)}$.
 Proof.
@@
 Let $f:=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$, where convergence is in $\mathrm{L}^{2}(\nu)$ at first.
-To show convergence in $H_{\nu}^{(\theta,r)}$-norm, we notice that $\ell^{(\mu,\theta,r)}\triangleq[\ell^{2}(I),\allowbreak\ell^{2}(\mu^{-1})]_{(\theta,r)}$ by Lemma 5.2.
+To show convergence in $H]_{\sim}^{(\theta,r)}-$norm, we notice that $\ell^{(\mu,\theta,r)}\triangleq[\ell^{2}(I),\allowbreak\ell^{2}(\mu^{-1})]_{(\theta,r)}$ by Lemma 5.2.
 Applying Lemma A.2 yields $\ell^{(\mu,\theta,r)}\hookrightarrow\ell^{2}(I)$.
 If $C$ denotes the corresponding operator norm, we thus find
-$\left\|f-\sum_{i=1}^{n}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{(\mu,\theta,r)}\leq C\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{\ell^{2}(I)}\xrightarrow{n\rightarrow\infty}0\text{.}$ (61)
+$\left\|f-\sum_{i=1}^{n}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{(\mu,\theta,r)}\leq C\|(b_{i}\mathbbm{1}_{i\geq n+1})\|_{\ell^{2}(I)}\xrightarrow{n\rightarrow\infty}0\text{.}$ (75)
 Liftings
@@
 $\mathcal{L}^{\infty}(X)\rightarrow\mathrm{L}^{\infty}(\nu),\quad f\mapsto[f]_{\sim}$ (76)
-which assigns to each $f$ the equivalence class of functions.
+which assigns to each $f$ its equivalence class of functions.
 This mapping is surjective by definition of $\mathrm{L}^{\infty}(\nu)$, hence it possesses a, not necessarily unique, right inverse.
@@
 for all $f\in\mathcal{L}^{\infty}(\nu)$.
-The proof is based on lifting theory as developed in [8]. For a derivation of the specific result, see [18, Equation (58)].
-So in other words, for each equivalence class in $\mathrm{L}^{\infty}(\nu)$, the operator $\phi$ chooses a bounded representative in a linear and continuous way.
-Auxiliary results for the proofs of Theorems 3.2 and 3.3
-Let us now introduce a setup which makes it possible to apply our results for direct sums of Banach spaces from Definition B.1.
+The proof is based on lifting theory as developed in [12]. For a derivation of the specific result, see [27, Equation (58)].
+In other words, for each equivalence class in $\mathrm{L}^{\infty}(\nu)$, the operator $\phi$ chooses a bounded representative in a linear and continuous way.
+Proofs of Theorems 3.2, 3.3, and 3.4
+Let us now introduce a setup which makes it possible to apply the results for direct sums of Banach spaces from Appendix B.
 Remark 5.8.
@@
 $\left|\sum_{j\in\mathbb{Z}}\left\langle\pi_{j}(a)\,{,}\,x_{j}\right\rangle_{E_{j}}\right|\leq C\|x\|_{\bigoplus_{j\in\mathbb{Z}}^{(r)}E_{j}}$ (90)
-where we replaced $\pi^{(r)}(b)$ by $x$ and applied the isometry property on the right hand side.
+where we replaced $\pi^{(r)}(b)$ by $x$ and applied the isometry property on the right-hand side.
 An application of Proposition B.5 shows that this is the case if and only if $(\pi_{j}(a))_{j\in\mathbb{Z}}\in\bigoplus_{j\in\mathbb{Z}}^{(r^{\prime})}E_{j}$ with norm $\|(\pi_{j}(a))_{j\in\mathbb{Z}}\|_{r^{\prime}}\leq C$.
@@
 ∎
+In order to prove Theorem 3.2, we will use liftings from Theorem 5.7. As this requires a complete measure space, which is not necessarily given in Theorem 3.2, we consider the following lemma.
+Lemma 5.11.
+[…]
+  1. (i)
+    For every $\hat{\mathcal{A}}-$measurable $f:X\rightarrow\mathbb{R}$, there exists a $\mathcal{\nu}$-null set $N$ such that $\mathbbm{1}_{X\setminus N}f$ is $\mathcal{A}$-measurable.
+  2. (ii)
+    The operator
+    $\mathrm{L}^{\infty}(\nu)\rightarrow\mathrm{L}^{\infty}(\hat{\nu}),\quad[f]_{\sim}\mapsto[f]_{\hat{\sim}}$ (91)
+    is a well-defined isometric isomorphism, where $\cdot\,$$_{\hat{\sim}}$ denotes equivalence classes with respect to $\hat{\nu}$.
+Proof.
+(i).
+By [6, Proposition 2.2.5], there are $\mathcal{A}$-measurable functions $f_{0}$ and $f_{1}$ on $X$ with $f_{0}\leq f\leq f_{1}$ and $f_{0}=f_{1}$ $\nu$-almost everywhere.
+Choosing $N=\{x\in X\,|\,f_{0}(x)\neq f_{1}(x)\}$ shows that $\mathbbm{1}_{X\setminus N}f=\mathbbm{1}_{X\setminus N}f_{0}$ is $\mathcal{A}$-measurable.
+(ii).
+We first prove that the mapping is well-defined.
+As $\mathcal{A}\subset\hat{\mathcal{A}}$, we find that any $\mathcal{A}$-measurable $f:X\rightarrow\mathbb{R}$ is $\hat{\mathcal{A}}-$measurable, which justifies writing $f]_{\hat{\sim}}$.
+To see that the choice of representative of $f]_{\sim}$ is irrelevant, let $f$ and $g$ be $\mathcal{A}$-measurable functions with $f]_{\sim}=[g]_{\sim}$.
+Then $f-g]_{\sim}=0$, hence there is a $\nu$-null set $N$ such that f(x)g(x)=0f(x)-g(x)=0 for all $x\in X\setminus N$.
+But by definition of the completion, $N$ is a $\hat{\nu}-$zero set as well, which shows that $f-g]_{\hat{\sim}}=0$.
+Hence we get $f]_{\hat{\sim}}=[g]_{\hat{\sim}}$.
+The linearity of the mapping is obvious. Moreover, for an $\mathcal{A}$-measurable $f:X\rightarrow\mathbb{R}$ with $f]_{\sim}\in\mathrm{L}^{\infty}(\nu)$, we have
+$\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$ $=\inf\{B\geq 0\,|\,\nu(\{x\in X\,|\,f(x)>B\})=0\}$ (92)
+$=\inf\{B\geq 0\,|\,\hat{\nu}(\{x\in X\,|\,f(x)>B\})=0\}$ (93)
+$=\|[f]_{\hat{\sim}}\|_{\mathrm{L}^{\infty}(\hat{\nu})}\text{,}$ (94)
+[…]
 The following lemma gives some details on the construction of $\overline{H}_{\nu}^{(\theta,r)}$ in Theorem 3.2.
 Lemma 5.12.
-Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+. Let
-$\overline{e}_{i}:=\phi([e_{i}]_{\sim})\text{,}$ (77)
-where $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ is the lifting according to Theorem 5.7.
-Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous.
-Then the following statements are true.
+Let $(X,\mathcal{A},\nu)$, $H$, $\mu_{i}$, and $e_{i}$ be as in Assumption H+.
+Moreover, let $0<\theta<1$ and $1\leq r\leq\infty$ be such that the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$ exists and is continuous.
+Then for all $i\in I$, there exist representatives $\overline{e}_{i}\in\mathcal{L}^{\infty}(X)$ of $e_{i}]_{\sim}$ such that the following statements are true:
   1. (i)
@@
   2. (ii)
-    It holds that $\sum_{i\in I}b_{i}\overline{e}_{i}=\sum_{i\in I}b_{i}e_{i}$ almost everywhere for all $b\in\ell^{(\mu,\theta,r)}$.
+    There is a $\nu$-null set $N\subset X$ such that $\sum_{i\in I}b_{i}\overline{e}_{i}(x)=\sum_{i\in I}b_{i}e_{i}(x)$ for all $b\in\ell^{(\mu,\theta,r)}$ and all $x\in X\setminus N$.
   3. (iii)
@@
     The mapping
-    $H_{\nu}^{(\theta,r)}\rightarrow\overline{H}_{\nu}^{(\theta,r)},\quad\sum_{i\in I}b_{i}[e_{i}]_{\sim}\mapsto\sum_{i\in I}b_{i}\overline{e}_{i}$ (79)
+    $[H]_{\sim}^{(\theta,r)}\rightarrow\overline{H}_{\nu}^{(\theta,r)},\quad\sum_{i\in I}b_{i}[e_{i}]_{\sim}\mapsto\sum_{i\in I}b_{i}\overline{e}_{i}$ (96)
     is an isometric isomorphism.
@@
     It holds that
-    $\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (80)
+    $\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (97)
 Proof.
-As the RKHS $H$ has a bounded kernel, the eigenfunctions $e_{i}$ are bounded as well.
-Hence by Theorem 5.7, the lifted eigenfunctions $\overline{e}_{i}=\phi([e_{i}]_{\sim})$ are measurable, bounded and we know that for each $i\in I$, there is a $\nu$-null set $N_{i}$ such that $e_{i}(x)=\overline{e}_{i}(x)$ for all $x\in X\setminus N_{i}$.
-As $I$ is at most countable, $N:=\bigcup_{i\in I}N_{i}$ is again measurable with $\nu(N)=0$.
-By that, for all $x\in X\setminus N$ and all $i\in I$, we have $\overline{e}_{i}(x)=e_{i}(x)$.
-Furthermore, let us denote
-$\kappa:=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (81)
+[…]
+Let us consider the $\nu$-null set $N:=N^{\prime}\cup\bigcup_{i\in I}N^{\prime\prime}_{i}$.
+Then for all $i\in I$, it holds that $\overline{e}_{i}:=\mathbbm{1}_{X\setminus N}\breve{e}_{i}$ is $\mathcal{A}$-measurable with $\overline{e}_{i}(x)=e_{i}(x)$ for all $x\in X\setminus N$.
+Let us denote the linear operator
+[…]
+which is well-defined by Lemma 5.11 and has operator norm at most $1$ due to Lemma 5.11 and Theorem 5.7. Note that $A[e_{i}]_{\sim}=\overline{e}_{i}$ are $\mathcal{A}$-measurable functions on $X$.
+Furthermore, we write
+$\kappa:=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (99)
 (i).
 Let $b\in\ell^{(\mu,\theta,r)}$.
-Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H_{\nu}^{(\theta,r)}$.
-Due to the continuity of the embedding $H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it converges in $\mathrm{L}^{\infty}(\nu)$ as well and we have
-$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{H_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (82)
-By continuity of the lifting operator $\phi:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$ from Theorem 5.7, we thus find
-$\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)=\sum_{i\in I}b_{i}\phi([e_{i}]_{\sim})=\sum_{i\in I}b_{i}\overline{e}_{i}\text{,}$ (83)
+Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H]_{\sim}^{(\theta,r)}$.
+Due to the continuity of the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it also converges in $\mathrm{L}^{\infty}(\nu)$, and we have
+$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (100)
+By continuity of the operator $A:\mathrm{L}^{\infty}(\nu)\rightarrow\mathcal{L}^{\infty}(X)$, we thus find
+$A\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)=\sum_{i\in I}b_{i}A([e_{i}]_{\sim})=\sum_{i\in I}b_{i}\overline{e}_{i}\text{,}$ (101)
 so $\sum_{i\in I}b_{i}\overline{e}_{i}$ converges in $\mathcal{L}^{\infty}(X)$, that is uniformly.
@@
 To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$.
-Because $\phi$ has operator norm at most $1$, by (83) we find that
-$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|\phi\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)\right\|_{\mathcal{L}^{\infty}(X)}\leq\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\text{.}$ (84)
-Now for all $x\in X$, we consider $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ to find that
-$\left|\delta_{x}\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)\right|=\left|\sum_{i\in I}b_{i}\overline{e}_{i}(x)\right|\leq\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}$ (85)
-where for second to last inequality, we combined (82) and (84).
-(iv). This follows as the norms on $H_{\nu}^{(\theta,r)}$ and on $\overline{H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
-(v). Note that by using (85), we find that
-$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}=\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{H_{\nu}^{(\theta,r)}}\text{,}$ (86)
-which shows $\leq$. For $\geq$, we note that
-$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|f\|_{\mathcal{L}^{\infty}(X)}\geq\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|H_{\nu}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|$ (87)
-where the last step follows as $\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}=\|[f]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$. ∎
+Because $A$ has operator norm at most $1$, we find by (101) that
+$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|A\left(\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right)\right\|_{\mathcal{L}^{\infty}(X)}\leq\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\text{.}$ (102)
+Now let $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ be the evaluation functional at some fixed $x\in X$. Combining (100) and (102), we find
+$\left|\delta_{x}\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)\right|=\left|\left(\sum_{i\in I}b_{i}\overline{e}_{i}\right)(x)\right|\leq\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (103)
+(iv). This follows as the norms on $H]_{\sim}^{(\theta,r)}$ and on $\overline{H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
+(v). Note that by using (100) and (102), we find
+$\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\mathcal{L}^{\infty}(X)}\leq\kappa\|b\|_{(\mu,\theta,r)}=\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{[H]_{\sim}^{(\theta,r)}}\text{,}$ (104)
+which shows $\leq$. For $\geq$, we use
+$\|\overline{H}_{\nu}^{(\theta,r)}\hookrightarrow\mathcal{L}^{\infty}(X)\|=\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|f\|_{\mathcal{L}^{\infty}(X)}\geq\sup_{\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}\leq 1}\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|$ (105)
+where the last step follows by $\|f\|_{\overline{H}_{\nu}^{(\theta,r)}}=\|[f]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}$. ∎
 Proof of Theorem 3.2.
-The assertion follows directly from Lemma 5.11 by using the inverse of the isometric isomorphism from (iv). ∎
+The assertion follows directly from Lemma 5.12 by using the inverse of the isometric isomorphism from (iv). The additional statement follows from (ii). ∎
+In Theorem 3.2, we describe by means of lifting theory how to find representatives of $H]_{\sim}^{(\theta,r)}$ that give rise to a Banach space of functions.
+However, note that we already chose a set of “canonical” representatives in Definition 2.3.
+Let us analyse under which circumstances this set of “canonical” representatives forms a BSF.
+To this end, let us apply a topological technique similar to [23].
 Proof of Theorem 3.3.
-First of all, note that by Definition 2.2, an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ can be written as a series
+Let us denote
+$\kappa:=\|[H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)\|\text{.}$ (106)
+(i).
+Let $b\in\ell^{(\mu,\theta,r)}$.
+Then by Lemma 5.6, we know that $\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ converges in $H]_{\sim}^{(\theta,r)}$.
+Due to the continuity of the embedding $H]_{\sim}^{(\theta,r)}\hookrightarrow\mathrm{L}^{\infty}(\nu)$, it converges in $\mathrm{L}^{\infty}(\nu)$ as well, and we have
+$\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{[H]_{\sim}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (107)
+Now the partial sums of $\sum_{i\in I}b_{i}e_{i}$ are $\tau(H)$-continuous and $\sum_{i\in I}b_{i}[e_{i}]_{\sim}=[\sum_{i\in I}b_{i}e_{i}]_{\sim}$ converges in $\mathrm{L}^{\infty}(\nu)$-norm.
+By Lemma C.4, we hence find that $\sum_{i\in I}b_{i}e_{i}$ converges in $C_{b}(X,\tau(H))$, hence uniformly.
+(ii). First of all, note that the coefficients $b_{i}$ are uniquely determined by the function $\sum_{i\in I}b_{i}e_{i}$, as the $e_{i}]_{\sim}$ are $\mathrm{L}^{2}(\nu)$-orthogonal.
+As $\ell^{(\mu,\theta,r)}$ is a Banach space by Remark 5.8, it follows that ${H}_{\nu}^{(\theta,r)}$ is a Banach space as well.
+By the uniform convergence in (i), ${H}_{\nu}^{(\theta,r)}$ consists of functions.
+It remains to show that the evaluation functionals are continuous.
+To this end, let us fix a $b\in\ell^{(\mu,\theta,r)}$.
+Then for all $x\in X$, we consider $\delta_{x}:\mathcal{L}^{\infty}(X)\rightarrow\mathbb{R},f\mapsto f(x)$ to find that
+$\begin{split}\left|\delta_{x}\left(\sum_{i\in I}b_{i}e_{i}\right)\right|=\left|\left(\sum_{i\in I}b_{i}e_{i}\right)(x)\right|\leq\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{\mathcal{L}^{\infty}(X)}=\left\|\sum_{i\in I}b_{i}[e_{i}]_{\sim}\right\|_{\mathrm{L}^{\infty}(\nu)}\leq\kappa\|b\|_{(\mu,\theta,r)}\text{,}\end{split}$ (108)
+where for second to last step, we used Corollary C.3, and the last inequality is due to (107).
+(iii). This follows as the norms on $H]_{\sim}^{(\theta,r)}$ and on ${H}_{\nu}^{(\theta,r)}$ are both inherited from the $\ell^{(\mu,\theta,r)}$-norm of the unique Fourier coefficients.
+(iv). In the proof of (i), we found that ${H}_{\nu}^{(\theta,r)}$ consists of $\tau(H)$-continuous and bounded functions.
+Note that by using the isometric isomorphism $\alpha:{H}_{\nu}^{(\theta,r)}\rightarrow[H]_{\sim}^{(\theta,r)}$ from (iii) and the isometry $\beta:\mathrm{C}_{b}(X,\tau(H))\rightarrow\mathrm{L}^{\infty}(\nu),f\mapsto[f]_{\sim}$, the following diagram commutes:
+${H}_{\nu}^{(\theta,r)}$$\mathrm{C}_{b}(X,\tau(H))$$H]_{\sim}^{(\theta,r)}$$\mathrm{L}^{\infty}(\nu)$$\scriptstyle{\mathrm{Id}}\scriptstyle{\simeq}$$\scriptstyle{\alpha}$$\scriptstyle{\beta}$$\scriptstyle{\mathrm{Id}}$
+But as $\mathrm{Id}:[H]_{\sim}^{(\theta,r)}\rightarrow\mathrm{L}^{\infty}(\nu)$ is bounded with operator norm $\kappa$, it follows that $\mathrm{Id}:{H}_{\nu}^{(\theta,r)}\rightarrow\mathrm{C}_{b}(X,\tau(H))$ is bounded with operator norm $\kappa$ as well.
+(v). Let $f\in H$. Then by Lemma 5.3, there is a sequence $(b_{i})_{i\in I}\in\ell^{2}(\mu^{-1})$ such that $f]_{\sim}=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$. It holds that
+$\ell^{2}(\mu^{-1})\hookrightarrow[\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}\triangleq\ell^{(\mu,\theta,r)}\text{,}$ (109)
+where for the first step, we apply Lemma A.2 and for the second, we apply Lemma 5.2 to the counting measure on $I$ and the weight sequence $\mu^{-1}$ to find that the spaces $\ell^{2}(I),\ell^{2}(\mu^{-1})]_{(\theta,r)}$ and $\ell^{(\mu,\theta,r)}$ are norm equivalent.
+We hence find that $(b_{i})_{i\in I}\in\ell^{(\mu,\theta,r)}$.
+Now by (i), the sum $\sum_{i\in I}b_{i}e_{i}$ converges uniformly in $C_{b}(X,\tau(H))$.
+Because $f$ itself is $\tau(H)$-continuous as well, we have $f=\sum_{i\in I}b_{i}e_{i}$, as almost everywhere agreeing $\tau(H)$-continuous functions agree everywhere by Corollary C.3.
+This shows that $f=\sum_{i\in I}b_{i}e_{i}\in H^{(\theta,r)}$.
+∎
+Proof of Theorem 3.4.
+First of all, note that by Definition 2.3, an arbitrary $g]_{\sim}\in[H]_{\sim}^{(\theta,r)}$ can be written as a series
 $[g]_{\sim}=\sum_{i\in I}b_{i}[e_{i}]_{\sim}$ (110)
-for a suitable $b\in\ell^{(\mu,\theta,r)}$, where the convergence is in $H_{\nu}^{(\theta,r)}$ by Lemma 5.6. In addition, it holds that
-$\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}=\|b\|_{(\mu,\theta,r)}\text{.}$ (89)
+for a suitable $b\in\ell^{(\mu,\theta,r)}$, where the convergence is in $H]_{\sim}^{(\theta,r)}$ by Lemma 5.6. In addition, it holds that
+$\|[g]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}=\|b\|_{(\mu,\theta,r)}\text{.}$ (111)
 For the rest of this proof, we write $f_{i}(x):=e_{i}(x)\mu_{i}^{\theta}$ for all $i\in I$ and all $x\in X$.
-“(i) $\Rightarrow$ (ii)”: By Property (i), there is a $\nu$-null set $N$ such that
+Concerning the different assumptions, note that Assumptions H+ and T imply Assumption H.
+“(i) $\Rightarrow$ (ii)” Obvious.
+“(ii) $\Rightarrow$ (iii)” By Property (ii), there is a $\nu$-null set $N$ such that
 $\left\|\left(f_{i}(x)\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}\leq\kappa$ (112)
 holds for all $x\in X\setminus N$. Hence it follows for $x\in X\setminus N$ and all $b\in\ell^{(\mu,\theta,r)}$ that
-$\sum_{i\in I}|b_{i}e_{i}(x)|$ $=\sum_{i\in I}|b_{i}f_{i}(x)|\mu_{i}^{-\theta}$ (91)
-$\leq\|(f_{i}(x))_{i\in I}\|_{(\mu,\theta,r^{\prime})}\|b\|_{(\mu,\theta,r)}$ (92)
-$\leq\kappa\,\|b\|_{(\mu,\theta,r)}\text{,}$ (93)
+$\begin{split}\sum_{i\in I}|b_{i}e_{i}(x)|&=\sum_{i\in I}|b_{i}f_{i}(x)|\mu_{i}^{-\theta}\\ &\leq\|(f_{i}(x))_{i\in I}\|_{(\mu,\theta,r^{\prime})}\|b\|_{(\mu,\theta,r)}\\ &\leq\kappa\,\|b\|_{(\mu,\theta,r)}\text{,}\end{split}$ (113)
 […]
-Consequently, we find for such an arbitrary $g]_{\sim}\in H_{\nu}^{(\theta,r)}$ that
 $\|[g]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$ $=\inf\left\{B\geq 0\,:\,\exists M\text{ with }\nu(M)=0\text{ and }\sup_{x\in X\setminus M}\left|g(x)\right|\leq B\right\}$ (114)
@@
 $\leq\kappa\|b\|_{(\mu,\theta,r)}$ (117)
-$=\kappa\|[g]_{\sim}\|_{H_{\nu}^{(\theta,r)}}$ (98)
-by an application of (93). This proves (ii).
-“(ii) $\Rightarrow$ (i)”: By Lemma 5.11, for almost all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
-$\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}=\sum_{i\in I}b_{i}e_{i}(x)=\sum_{i\in I}b_{i}\overline{e}_{i}(x)\leq\kappa\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\overline{H}_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (99)
-Thus, for almost all $x\in X$, an application of Lemma 5.10 to $a_{i}=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$.
-This finally shows (i).
+$=\kappa\|[g]_{\sim}\|_{[H]_{\sim}^{(\theta,r)}}$ (118)
+by an application of (113). This proves (iii).
+“(iii) $\Rightarrow$ (ii)” under Assumption H+: By Lemma 5.12, for almost all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
+$\left|\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}\right|=\left|\sum_{i\in I}b_{i}e_{i}(x)\right|=\left|\sum_{i\in I}b_{i}\overline{e}_{i}(x)\right|\leq\kappa\left\|\sum_{i\in I}b_{i}\overline{e}_{i}\right\|_{\overline{H}_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{, }$ (119)
+where $\overline{e}_{i}$ is defined as in Lemma 5.12.
+Note that the sums are absolutely convergent by Remark B.2.
+For almost all $x\in X$, an application of Lemma 5.10 to $a_{i}:=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$.
+This shows (ii).
+“(ii) $\Rightarrow$ (i)” under Assumption T: An application of Lemma C.1 shows that the function $X\rightarrow[0,\infty]$ given by
+$x\mapsto\left\|\left(e_{i}(x)\mu_{i}^{\theta}\right)_{i\in I}\right\|_{(\mu,\theta,r^{\prime})}$ (120)
+is non-negative and lower semicontinuous. Hence by Lemma C.2, the essential supremum and the supremum agree.
+“(iii) $\Rightarrow$ (ii)” under Assumption T: By Theorem 3.3, for all $x\in X$ and all $b\in\ell^{(\mu,\theta,r)}$, we have
+$\left|\sum_{i\in I}b_{i}f_{i}(x)\mu_{i}^{-\theta}\right|=\left|\sum_{i\in I}b_{i}e_{i}(x)\right|\leq\kappa\left\|\sum_{i\in I}b_{i}e_{i}\right\|_{H_{\nu}^{(\theta,r)}}=\kappa\|b\|_{(\mu,\theta,r)}\text{.}$ (121)
+Thus, for all $x\in X$, another application of Lemma 5.10 to $a_{i}:=f_{i}(x)$ yields that $(f_{i}(x))_{i\in I}\in\ell^{(\mu,\theta,r^{\prime})}$ with norm $\|f_{i}(x)\|_{(\mu,\theta,r^{\prime})}\leq\kappa$.
+This shows (ii).
 ∎
@@
 with convergence in $\mathrm{L}^{2}(P_{X})$.
-Then by the norm equivalence from Theorem 3.1, we find a $C_{6}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
-$\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{6}\|(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}\|_{(\mu,\theta,r)}$ (110)
-$\leq C_{6}\sup_{i\in I}\left(\big{|}1-\mu_{i}g_{\lambda}(\mu_{i})\big{|}\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (111)
+Then by the norm equivalence from Theorem 3.1, we find a $C_{5}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.13:
+$\|[f_{P,\lambda,g}]_{\sim}-f_{L,P}^{*}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{5}\|(1-\mu_{i}g_{\lambda}(\mu_{i}))b_{i}\|_{(\mu,\theta,r)}$ (132)
+$\leq C_{5}\sup_{i\in I}\left(\big|1-\mu_{i}g_{\lambda}(\mu_{i})\big|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (133)
 From this, the assertion follows by the upper bound on $\|b\|_{(\mu,\beta,r)}$ given in (128).
-(ii). By the norm equivalence in Theorem 3.1, we find a $C_{5}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.12:
-$\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{5}\|\mu_{i}g_{\lambda}(\mu_{i})b_{i}\|_{(\mu,\theta,r)}$ (112)
-$\leq C_{5}\sup_{i\in I}\left(|\mu_{i}g_{\lambda}(\mu_{i})|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (113)
+(ii). By the norm equivalence in Theorem 3.1, we find a $C_{6}\geq 0$ independent of $f_{L,P}^{*}$ such that we can estimate as follows via Lemma 5.13:
+$\|[f_{P,\lambda,g}]_{\sim}\|_{[\mathrm{L}^{2}(X),[H]_{\sim}]_{\theta,r}}$ $\leq C_{6}\|\mu_{i}g_{\lambda}(\mu_{i})b_{i}\|_{(\mu,\theta,r)}$ (134)
+$\leq C_{6}\sup_{i\in I}\left(|\mu_{i}g_{\lambda}(\mu_{i})|\mu_{i}^{\frac{\beta-\theta}{2}}\right)\|b\|_{(\mu,\beta,r)}$ (135)
 From this, the estimate follows by the upper bound on $\|b\|_{(\mu,\beta,r)}$ given in (128). ∎
@@
 We begin by proving that isometric isomorphisms can be carried over from Banach spaces $E_{2}\hookrightarrow E_{1}$ to the interpolation spaces $E_{1},E_{2}]_{\theta,r}$.
-For a map $f:A\rightarrow B$ and $A^{\prime}\subset A$, $B^{\prime}\subset B$, we use the notation $f|_{A^{\prime}\rightarrow B^{\prime}}$ for the map $A^{\prime}\rightarrow B^{\prime}$ given by $x\mapsto f(x)$ in case it is well defined.
+For a map $f:A\rightarrow B$ and $A^{\prime}\subset A$, $B^{\prime}\subset B$, we use the notation $f|_{A^{\prime}\rightarrow B^{\prime}}$ for the map $A^{\prime}\rightarrow B^{\prime}$ given by $x\mapsto f(x)$ in case it is well-defined.
 Lemma A.1.
 Let $E_{2}\hookrightarrow E_{1}$ and $F_{2}\hookrightarrow F_{1}$ be Banach spaces.
-Suppose there is a bounded linear operator $\phi:E_{1}\rightarrow F_{1}$ such that $\phi|_{E_{1}\rightarrow F_{1}}$ and $\phi|_{E_{2}\rightarrow F_{2}}$ are isometric isomorphisms.
+Suppose there is a bounded linear operator $\phi:E_{1}\rightarrow F_{1}$ such that $\phi$ and $\phi|_{E_{2}\rightarrow F_{2}}$ are isometric isomorphisms.
 Then for $0<\theta<1$ and $1\leq r\leq\infty$, the following map is an isometric isomorphism:
@@
 Our first goal is to show
-$[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}\subset H_{1}$ (120)
+$[H,H_{2}]_{\theta,r}\cup[H_{1},H_{2}]_{\theta,r}\subset H_{1}\text{.}$ (142)
 To this end, we first note that $H_{1},H_{2}]_{\theta,r}\subset H_{1}$ by Lemma A.2.
@@
 B Appendix: Direct sums of Banach spaces
-A closer look at Definition 2.1 reveals that the $\ell^{(\mu,\theta,r)}$-norms are built up from two nested norms: The inner part (10) of the $(\mu,\theta,r)$-norm originates from an $\ell^{2}(\mu^{-\theta})$-norm taken over the restricted indices in $M_{j}$, while the outer sum in (11) corresponds to the $\ell^{r}$-norm taken over the values given by the inner part.
+A closer look at Definition 2.2 reveals that the $\ell^{(\mu,\theta,r)}$-norms are built up from two nested norms: The inner part (10) of the $(\mu,\theta,r)$-norm originates from an $\ell^{2}(\mu^{-\theta})$-norm taken over the restricted indices in $M_{j}$, while the outer part in (11) corresponds to the $\ell^{r}$-norm taken over the values given by the inner part.
 To deal with these norms, let us introduce such nested norms in a more general setting. For notational convenience, let us assume that $I=\{1,...,n\}$ or $I=\mathbb{N}$ is an index set.
@@
 Then we find that $x_{i}^{\prime}(\tilde{x}_{i})=|x_{i}^{\prime}(x_{i})|$, and we still have $\|\tilde{x}\|_{p}=\|x\|_{p}$.
-Hence using $\tilde{x}$ instead of $x$ on the left hand side gives the right hand side.
+Hence using $\tilde{x}$ instead of $x$ on the left-hand side gives the right-hand side.
 In addition, note that due to
@@
 ${\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}\rightarrow\left({\bigoplus_{i\in I}}^{(p)}E_{i}\right)^{\prime},\quad x^{\prime}\mapsto\left(x\mapsto\sum_{i\in I}x^{\prime}_{i}(x_{i})\right)$ (159)
-is an
+is
   1. (i)
-    isometric isomorphism if $p=0$ and $p^{\prime}=1$.
+    an isometric isomorphism if $p=0$ and $p^{\prime}=1$,
   2. (ii)
-    isometric isomorphism if $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$.
+    an isometric isomorphism if $1\leq p<\infty$ with $\frac{1}{p}+\frac{1}{p^{\prime}}=1$, and
   3. (iii)
-    isometry if $p=\infty$ and $p^{\prime}=1$.
+    an isometry if $p=\infty$ and $p^{\prime}=1$.
 Proof.
@@
 Let us prove (iii) here.
-We begin by showing that the operator is well defined in the sense that the sum converges.
+We begin by showing that the operator is well-defined in the sense that the sum converges.
 In fact, we even have absolute convergence, as for $x^{\prime}\in{\bigoplus_{i\in I}}^{(1)}E_{i}^{\prime}$ and $x\in{\bigoplus_{i\in I}}^{(\infty)}E_{i}$, we find that
@@
 $\sup_{n\in\mathbb{N}}|S_{n}(x)|=\sup_{n\in\mathbb{N}}\left|\sum_{i=1}^{n}x^{\prime}_{i}(x_{i})\right|\leq\sup_{n\in\mathbb{N}}\sum_{i=1}^{n}|x^{\prime}_{i}(x_{i})|\leq\sum_{i=1}^{\infty}|x^{\prime}_{i}(x_{i})|<\infty\text{,}$ (170)
-where we used the absolute convergence of (145). Hence the uniform boundedness principle [13, Theorem 4.52] yields that $\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$.
+where we used the absolute convergence of (167). Hence the uniform boundedness principle [21, Theorem 4.52] yields that $\sup_{n\in\mathbb{N}}\|S_{n}\|<\infty$.
 Our next goal is to show
@@
 To this end, note that for the operator norm we find the following:
-$\|S_{n}\|=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}((x_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\leq n}^{(0)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}\text{.}$ (150)
+$\|S_{n}\|=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\in\mathbb{N}}^{(\infty)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}((x_{i}\mathbbm{1}_{i\leq n})_{i\in\mathbb{N}}|}{\|x\|_{\infty}}=\sup_{\begin{subarray}{c}x\in\bigoplus_{i\leq n}^{(0)}E_{i}\\ x\neq 0\end{subarray}}\frac{|S_{n}x|}{\|x\|_{\infty}}$ (172)
 This means that the operator norm of $S_{n}$ equals the operator norm of $S_{n}$ restricted to ${\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i}$.
-This restricted operator is as an element of $({\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i})^{\prime}\simeq{\bigoplus_{i\in\mathbb{N}}}^{(1)}E_{i}^{\prime}$, where we are using the isometric isomorphism from Lemma B.3.
+This restricted operator is an element of $({\bigoplus_{i\in\mathbb{N}}}^{(0)}E_{i})^{\prime}\simeq{\bigoplus_{i\in\mathbb{N}}}^{(1)}E_{i}^{\prime}$, where we are using the isometric isomorphism from Lemma B.3.
 This isometric isomorphism assigns $S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}$ to the sequence $(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\in\bigoplus_{i\leq n}^{(1)}E_{i}^{\prime}$.
 We hence find the equality of norms
-$\|S_{n}\|=\left\|S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}\right\|=\left\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\right\|_{1}\text{.}$ (151)
+$\|S_{n}\|=\left\|S_{n}|_{\bigoplus_{i\leq n}^{(0)}E_{i}}\right\|=\left\|(x^{\prime}_{i}\mathbbm{1}_{i\leq n})_{i\in I}\right\|_{1}$ (173)
 which finishes the second case.
@@
   2. (ii)
-    It holds that $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ with norm $\|x^{\prime}\|_{p^{\prime}}\leq C$.
+    It holds that $x^{\prime}\in{\bigoplus_{i\in I}}^{(p^{\prime})}E_{i}^{\prime}$ with $\|x^{\prime}\|_{p^{\prime}}\leq C$.
 Proof.
@@
 and note that the assertion follows by treating absolute values as explained in Remark B.2. ∎
+C Appendix: Semicontinuous functions
+Let $(X,\mathcal{T})$ be a topological space.
+Then a function $g:X\rightarrow[0,\infty]$ is called lower semicontinuous (l.s.c.) if and only if the sets $g^{-1}((a,\infty])$ are open in $X$ for all $a\geq 0$.
+Lemma C.1.
+Let $(X,\mathcal{T})$ be a topological space and suppose that $g_{i}:X\rightarrow[0,\infty]$ are l.s.c.
+functions for all $i$ in an index set $I$.
+Then the following functions $X\rightarrow[0,\infty]$ are l.s.c.:
+  1. (i)
+    $\sup_{i\in I}g_{i}$
+  2. (ii)
+    $\sum_{i\in I}g_{i}$
+  3. (iii)
+    $g_{1}\cdot g_{2}$
+  4. (iv)
+    $g_{1}^{\alpha}$ for all $\alpha>0$
+Proof.
+(i). Non-negativity follows directly, for lower semicontinuity, see [5, §6.2 Proposition 2].
+(ii).
+Non-negativity follows directly.
+By [5, §6.2 Proposition 2], finite sums of l.s.c.
+functions are l.s.c.
+The property follows for infinite sums as $\sum_{i\in I}g_{i}(x)=\sup_{j\in I}\sum_{i=1}^{j}g_{i}(x)$ holds for non-negative functions.
+(iii). Non-negativity follows directly, for lower semicontinuity, see [5, §6.2 Proposition 2].
+(iv). This follows from the fact that for all $a\geq 0$ we have
+$\{x\in X\,|\,g_{1}(x)^{\alpha}>a\}=\{x\in X\,|\,g_{1}(x)>a^{\frac{1}{\alpha}}\}\text{,}$
+which is an open set. ∎
+Lemma C.2.
+Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive.
+Then for all l.s.c.
+functions $g:X\rightarrow[0,\infty]$ it holds that
+$\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)=\sup_{x\in X}g(x)\text{.}$ (181)
+Proof.
+Assume the converse, that means $\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)<\sup_{x\in X}g(x)$. It follows that there are $\varepsilon>0$ and $y\in X$ such that the set
+$M:=\{y\in X\,|\,g(y)-\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)\in(\varepsilon,\infty]\}$ (182)
+is not empty.
+Additionally, $M$ is open as $g$ is l.s.c.
+By the positivity of $\nu$, we find that $\nu(M)>0$.
+Hence $g>\operatorname*{\mathrm{ess-sup}}_{x\in X}g(x)$ holds on the set of positive measure $M$, a contradiction.
+∎
+Corollary C.3.
+Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive.
+Let $f:X\rightarrow\mathbb{R}$ be continuous.
+Then $\|f\|_{\mathcal{L}^{\infty}(X)}=\|[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}$.
+Proof.
+Apply Lemma C.2 to the non-negative and $\mathcal{T}$-continuous function $|f|$. ∎
+Lemma C.4.
+Let $(X,\mathcal{T})$ be a topological space and $(X,\mathcal{A},\nu)$ be a measure space such that $\nu$ is $\mathcal{T}$-positive.
+Then $\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\}$ is a closed subspace of $\mathrm{L}^{\infty}(\nu)$ which is isometric isomorphic to $\mathrm{C}_{b}(X,\mathcal{T})$ by the mapping $f\mapsto[f]_{\sim}$ for $f\in\mathrm{C}_{b}(X,\mathcal{T})$.
+Proof.
+The fact that $\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\}$ is a linear subspace of $\mathrm{L}^{\infty}(\nu)$ follows from the definition.
+It remains to prove that it is closed.
+We do so by proving completeness, which is equivalent to closedness.
+To this end, let $(f_{n})_{n\in\mathbb{N}}$ be a sequence of functions in $\mathrm{C}_{b}(X,\mathcal{T})$ such that the $f_{n}]_{\sim}$ form a Cauchy sequence in $\mathrm{L}^{\infty}(\nu)$-norm.
+It follows by Corollary C.3 that for $n,m\rightarrow\infty$, we have
+$\|f_{n}-f_{m}\|_{\mathcal{L}^{\infty}(X)}=\|[f_{n}-f_{m}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[f_{n}]_{\sim}-[f_{m}]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}\rightarrow 0\text{.}$ (183)
+As $\mathrm{C}_{b}(X,\mathcal{T})$ with the $\mathcal{L}^{\infty}(X)$-norm is complete, there is a $f\in\mathrm{C}_{b}(X,\mathcal{T})$ such that $\|f_{n}-f\|_{\mathcal{L}^{\infty}(X)}\rightarrow 0$.
+But this implies
+$\|[f_{n}]_{\sim}-[f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|[f_{n}-f]_{\sim}\|_{\mathrm{L}^{\infty}(\nu)}=\|f_{n}-f\|_{\mathcal{L}^{\infty}(X)}\rightarrow 0$ (184)
+by another application of Corollary C.3. This shows that $f]_{\sim}$ is the desired limit of $([f_{n}]_{\sim})_{n\in\mathbb{N}}$ in $\mathrm{L}^{\infty}(\nu)$-norm.
+Finally, as the $\mathcal{L}^{\infty}(X)$-norm of a function $f\in\mathrm{C}_{b}(X,\tau(H))$ agrees with the $\mathrm{L}^{\infty}(\nu)$-norm of $f]_{\sim}$ by Corollary C.3, the map
+$\mathrm{C}_{b}(X,\mathcal{T})\rightarrow\{[f]_{\sim}\,|\,f\in\mathrm{C}_{b}(X,\mathcal{T})\},\quad f\mapsto[f]_{\sim}$ (185)
+is isometric, hence injective. Its surjectivity follows directly. ∎